### v1

# A Neurosymbolic Fast and Slow Architecture for Graph Coloring

Abstract
Constraint Satisfaction Problems (CSPs) present significant challenges to artificial intelligence due to their intricate constraints and the necessity for precise solutions. Existing symbolic solvers are often slow, and prior research has shown that Large Language Models (LLMs) alone struggle with CSPs because of their complexity. To bridge this gap, we build upon the existing SOFAI architecture (or SOFAI-v1), which adapts Daniel Kahneman’s “Thinking, Fast and Slow” cognitive model to AI. Our enhanced architecture, SOFAI-v2, integrates refined metacognitive governance mechanisms to improve adaptability across complex domains, specifically tailored here for solving CSPs like graph coloring. SOFAI-v2 combines a fast System 1 (S1) based on LLMs with a deliberative System 2 (S2) governed by a metacognition module. S1’s initial solutions, often limited by non-adherence to constraints, are enhanced through metacognitive governance, which provides targeted feedback and examples to adapt S1 to CSP requirements. If S1 fails to solve the problem, metacognition strategically invokes S2, ensuring accurate and reliable solutions. With empirical results, we show that SOFAI-v2 for graph coloring problems achieves an 16.98% increased success rate and is 32.42% faster than symbolic solvers.
Introduction
Constraint Satisfaction Problems are a core challenge in artificial intelligence (AI) due to their demand for correctness adhering to strict constraints in static (Kumar 1992) or uncertain, dynamic environments (Verfaillie and Jussien 2005). These problems are prominent in applications such as scheduling, boolean satisfiability, resource allocation, temporal reasoning, and planning (Welsh and Powell 1967; Chaitin 1982). Traditional approaches to solve CSPs typically rely on constraint propagation or search (Haralick and Elliott 1980). The symbolic approaches are known for their accuracy but often struggle with scalability as the complexity of CSPs increases, hindered by high computational overhead and slow processing times (Dechter 2003; Rossi, Van Beek, and Walsh 2006).
Recent advancements in LLMs have shown promise in rapidly processing complex information (Ruoss et al. 2024), but they are fundamentally limited in solving sequential decision-making tasks like planning (Valmeekam et al. 2022), and reasoning (Stechly, Valmeekam, and Kambhampati 2024). Due to their probabilistic, retrieval-based nature, LLMs approximate solutions rather than generating definitive answers, often leading to partial or inconsistent adherence to constraints (Jiang et al. 2023). Even with state-of-the-art prompting strategies and self-verification techniques, LLMs fall short in handling the strict requirements of CSPs, making their outputs unreliable for applications demanding high precision.
The disparity between the precision of traditional symbolic methods and the adaptability and speed of LLMs underscores the need for a new paradigm in tackling complex tasks like graph coloring (Jensen and Toft 2011), a canonical CSP. Drawing inspiration from Daniel Kahneman’s cognitive theory of “Thinking, Fast and Slow” (Kahneman 2011) and its adaptation to AI as SOFAI-v1 (Booch et al. 2021), we introduce SOFAI-v2 which extends the earlier architecture by refining metacognitive mechanisms that oversee a fast, experience-based S1, powered by an LLM, and a slow, deliberative S2 using the DSATUR algorithm for graph coloring. This governance mechanism continuously monitors and enhances S1’s outputs, offering targeted feedback and constraint adherence examples; when S1’s solutions fall short, S2 is strategically invoked to guarantee accuracy. In this setup, S1 swiftly produces initial solutions, while S2 provides a reliable fallback for achieving precise constraint satisfaction. Our contributions are as follows:
  • We introduce SOFAI-v2, a neurosymbolic fast and slow architecture designed to enhance the performance of CSP solvers on graph coloring problems through adaptive metacognitive governance.
  • We develop a comprehensive benchmark suite of graph coloring problems featuring variations in graph size, edge probability, and a mix of solvable and unsolvable instances to test solver adaptability.
  • We conduct extensive empirical evaluations of various solver configurations for graph coloring, demonstrating the advantages of SOFAI-v2 in terms of success rate, time efficiency, and adaptability.
The rest of the paper is organized as follows: we first provide the necessary background and discuss prior work. Next, we present the SOFAI-v2 architecture, followed by the experimental setup and results. Finally, we conclude with a summary of our findings.
Background
Constraint Satisfaction Problem
CSP is defined as a tuple $(X,D,C)$, where:
  • $X=\{x_{1},x_{2},\dots,x_{n}\}$ is a finite set of variables.
  • $D=\{D_{1},D_{2},\dots,D_{n}\}$ is a set of domains, with each $D_{i}$ representing the possible values for variable $x_{i}$.
  • $C=\{c_{1},c_{2},\dots,c_{m}\}$ is a set of constraints, with each constraint $c_{j}$ defining a subset of allowed assignments for a set of variables involved in the constraint.
A solution to a CSP is an assignment of values to all variables in $X$ such that every constraint in $C$ is satisfied. CSPs are used extensively in scheduling, resource allocation, and combinatorial optimization, where satisfying multiple constraints simultaneously is crucial.
Graph Coloring Problem
The graph coloring problem is a specific type of CSP, where the goal is to assign a color to each node of a graph such that no two adjacent nodes share the same color. Formally, it can be defined as a tuple $(G,C,k)$, where:
  • $G=(V,E)$ is an undirected graph, with $V=\{v_{1},v_{2},\dots,v_{n}\}$ representing the set of vertices (or nodes) and $E=\{(v_{i},v_{j})|v_{i},v_{j}\in V\}$ representing the set of edges.
  • $C=\{c_{1},c_{2},\dots,c_{k}\}$ is a set of $k$ colors.
  • The objective is to find a function $f:V\rightarrow C$ such that for any edge $(v_{i},v_{j})\in E$, $f(v_{i})\neq f(v_{j})$.
We use the DIMACS representation (Johnson and Trick 1996) to describe the graph coloring problem, and its corresponding visualization is shown in Figure 1.
1 c Example of a graph coloring problem in DIMACS format
2 p edge 5 5
3 c edges
4 e A B
5 e A C
6 e B C
7 e C D
8 e D E
(a) DIMACS Format of Graph Coloring Problem
(b) Visualization of the Graph
Figure 1: Graph Coloring Problem in DIMACS Format and Its Visualization
Degree of Saturation Algorithm with Backtracking
The Degree of Saturation, or DSATUR algorithm is a heuristic for approaching optimal coloring by assigning colors to vertices based on their saturation degree—the number of distinct colors already assigned to adjacent vertices (Brélaz 1979). While the algorithm by itself is not guaranteed to find an optimal solution, the addition of backtracking allows it to systematically explore alternative color assignments, ensuring optimal coloring is achieved.
Given the graph coloring problem $(G,C,k)$, where $G=(V,E)$ is an undirected graph, $C=\{c_{1},c_{2},\dots,c_{k}\}$ is a set of colors, and $f:V\rightarrow C$ is the coloring function, the algorithm proceeds as follows:
  1. 1.
    Initialization: Set $f(v)=\emptyset$ for all $v\in V$, marking all vertices as uncolored.
  2. 2.
    Saturation Degree Calculation: For each uncolored vertex $v\in V$, calculate its saturation degree, $d_{sat}(v)$, defined as the number of distinct colors assigned to its adjacent vertices: $\{f(u)\mid(v,u)\in E,f(u)\neq\emptyset\}$.
  3. 3.
    Vertex Selection: Choose the uncolored vertex $v$ with the highest saturation degree $d_{sat}(v)$. In the case of a tie, select the vertex with the highest degree $d(v)$, which is the number of edges incident to $v$.
  4. 4.
    Color Assignment: Assign the smallest feasible color $c\in C$ to $v$, ensuring that $f(v)\neq f(u)$ for all adjacent vertices $(v,u)\in E$.
  5. 5.
    Backtracking (to Ensure Optimality): If no valid color can be assigned without violating constraints, initiate backtracking. Backtracking systematically undoes recent assignments and explores alternate color choices. This recursive search allows the algorithm to revisit decisions, thereby exploring all possible color assignments and guaranteeing that an optimal solution—one using the minimal number of colors—is achieved.
  6. 6.
    Repeat: Continue steps 2-5 until all vertices are colored or no valid coloring exists.
The algorithm finds a minimal coloring assignment by prioritizing vertices with the highest constraints and employing backtracking to explore alternatives. Without backtracking, the heuristic alone may fall short of optimality; with backtracking, however, the algorithm can reach an optimal solution.
Related Works
Figure 2: SOFAI-v2 architecture with real-time metacognitive governance
This section reviews the evolution of solvers for graph coloring, the application of LLMs in sequential decision-making tasks, and the emergence of neurosymbolic approaches for enhancing decision-making processes.
Solvers for CSP
Traditional symbolic solvers have been extensively developed for this problem, leveraging algorithms that systematically explore the search space. Early works like Brélaz (1979) introduced efficient heuristics such as the DSATUR algorithm, significantly improving practical performance. Chaitin (1982) applied graph coloring to compiler optimization, demonstrating the versatility of symbolic methods. Golumbic (2004) provided a comprehensive treatment of graph algorithms, including coloring techniques. Despite their accuracy, symbolic approaches often struggle with scalability as CSP complexity increases, hindered by high computational overhead and slow processing times (Dechter 2003). Techniques like constraint propagation (Mackworth 1977) and backtracking algorithms (Bitner and Reingold 1975) have been employed to enhance efficiency but are limited by the exponential growth of the search space. To address these limitations, neural solvers have been explored, utilizing machine learning techniques to approximate solutions for graph coloring. For instance, Li, Chen, and Koltun (2018) proposed using graph neural networks to tackle combinatorial optimization problems. Similarly, Khalil et al. (2017) introduced a framework that learns heuristics directly from data, improving scalability. However, neural approaches often face challenges in strictly adhering to the constraints inherent in CSPs.
LLMs for Reasoning Tasks
Recent developments in LLMs have opened new avenues for addressing sequential decision-making tasks, including CSPs. Various prompt engineering techniques have been proposed to enhance the reasoning capabilities of LLMs (Brown 2020; Wei et al. 2022; Kojima et al. 2022; Nye et al. 2021; Wang et al. 2022). Han et al. (2023) demonstrated that LLMs can arrive at correct answers for graph-based problems through iterative prompting. However, the reasoning capabilities of LLMs have been found to be fundamentally approximate retrieval in nature (Kambhampati 2024), limiting their effectiveness in precise reasoning tasks. This limitation is evidenced by studies showing LLMs’ inability to solve planning problems (Valmeekam et al. 2022) and graph coloring problems (Stechly, Valmeekam, and Kambhampati 2024), where strict adherence to constraints is required (Jiang et al. 2023).
Neurosymbolic Approaches for Reasoning Tasks
Neurosymbolic approaches have gained considerable attention in addressing complex sequential decision-making tasks such as automated planning (Fabiano et al. 2023), constrained grid navigation (Ganapini et al. 2022), and puzzle solving (Lin et al. 2024). These studies demonstrate that integrating neural methods, such as LLMs, with symbolic solvers enhances performance on sequential decision-making tasks. To the best of our knowledge, neurosymbolic techniques have not been explored in the context of graph coloring problems.
The SOFAI Architecture
The SOFAI architecture, also referred to as SOFAI-v1 here (Booch et al. 2021), leverages a rule-based metacognitive control mechanism that dynamically chooses between S1 and S2 solvers based on a confidence threshold associated with S1’s outputs. S1 solvers are experience-based (and usually data-driven) solvers, while S2 solvers are deliberative (and usually symbolic and rule-based) solvers. SOFAI-v1 has been instantiated to planning (Fabiano et al. 2023) and constrained-grid navigation (Ganapini et al. 2022), showing improvement on the performance compared to either symbolic or data-driven solvers for the same class of problems.
In SOFAI-v1, the metacognitive component only chooses between S1 and S2 solvers, but does not exploit the possible collaboration between these two kinds of solvers, nor does it provide feedback to the solvers if they cannot solve the given problem instance. In this paper we consider a generalized version of this architecture, called SOFAI-v2, shown in Figure 2, where the metacognitive component (called MC) can provide feedback and examples to a failing S1 solver (that for this paper is an LLM), calling the S1 solver more than once, until either a correct solution is returned or a maximum number of iterations is reached, which triggers the activation of an S2 solver.
Formally, let $f_{\text{S1}}(x,M)$ denote the solution produced by S1 for a problem instance $x$ with access to an episodic memory $M$. Here, $M$ consists of previously encountered problem-solution pairs generated by SOFAI-v2, i.e., pairs $(x_{j},f_{\text{SOFAI-v2}}(x_{j}))$, where $f_{\text{SOFAI-v2}}(x_{j})$ represents the final solution that could have been generated by either S1 or S2.
When a new problem instance $x$ is presented to S1, a similarity function $\sigma(x,M)$ is used to retrieve a subset $M_{x}\subset M$ of similar past instances, defined as:
$M_{x}=\{(x_{j},f_{\text{SOFAI-v2}}(x_{j}))\mid\sigma(x,x_{j})\geq\alpha\},$
where $x_{j}\in M$ and $\alpha$ is a predefined similarity threshold. This subset $M_{x}$ provides S1 with additional contextual information from past solutions, enabling it to leverage the SOFAI-v2’s previous outputs for experience-based adaptation.
Let $C(f_{\text{S1}}(x,M_{x}))$ denote the confidence score assigned by MC to the solution $f_{\text{S1}}(x,M_{x})$, produced by S1 using both the current problem instance $x$ and the retrieved memory subset $M_{x}$. If $C(f_{\text{S1}}(x,M_{x}))\geq\theta_{\text{S1}}$, where $\theta_{\text{S1}}$ is a predefined confidence threshold, the solution is accepted as final. Otherwise, MC provides feedback $\mathcal{F}(f_{\text{S1}}(x,M_{x}))$ or generates a simplified example $\mathcal{E}(x)$, where $\mathcal{E}(x)$ is a reduced version of $x$, and calls the S1 solver again with the additional information included in the feedback and the examples.
MC also monitors the trend in S1’s confidence scores over $k$ iterations to evaluate the improvement. Let $C_{i}=C(f_{\text{S1}}(x_{i},M_{x_{i}}))$ represent the confidence score after the $i$-th iteration on similar or modified instances. An improvement trend is identified if:
$\forall i\in\{1,\dots,k-1\},\quad C_{i+1}>C_{i}.$
If this trend condition is not satisfied after $k$ iterations, MC gives the problem instance to S2, invoking $f_{\text{S2}}(x)$ as a fallback to solve it. This ensures robustness, particularly in cases where S1’s adaptation is insufficient to meet confidence requirements, even with feedback, examples, and episodic memory support.
Experimental Setup and Results
This section details the methodology and results of experiments designed to evaluate the performance of various solver configurations on graph coloring problems. The experiments focus on efficiency and success rate, incorporating a systematic approach to data generation, problem classification, and solver configuration.
Data Generation and Problem Classification
Graph coloring problems were generated using the Erdős–Rényi model, allowing for controlled variations in graph complexity through two main parameters:
Graph Size ($n$): Graphs varied $n\in[5,50]$ vertices, addressing a broad range of problem sizes.
Edge Formation Probability ($p$): Set with $p\in[0.1,0.9]$, this parameter adjusted the density of edges, thereby influencing the difficulty of the coloring problem.
Solvability Mix ($m$): The chromatic number for each graph, crucial for defining solvability, was calculated using the DSATUR algorithm. Based on this, problems were categorized into different solvability mixes, represented by the tuple $m=(a,b)$, where $a$ is the percentage of solvable problems and $b$ is the percentage of unsolvable problems:
  • $m=(100,0)$: All problems are solvable within the given constraints.
  • $m=(50,50)$: A balanced mix with equal solvable and unsolvable scenarios.
  • $m=(0,100)$: All problems are unsolvable, with the allowed number of colors less than the chromatic number.
Graphs were saved in DIMACS format to ensure standardization and enhance reproducibility.
SOFAI-v2 Implementation for Graph Coloring
The SOFAI-v2 architecture integrates advanced components designed to solve graph coloring problems dynamically. Figures 8 and 4 show example prompts provided to LLM (S1) for solving graph coloring. Below, we detail the implementation of these components using formal definitions and the functions they perform within the architecture:
  • Episodic Memory Implementation: The episodic memory, denoted as $M$, stores historical instances of solved graph coloring problems and their solutions. This aids the S1 solver by providing access to a set of similar past problems, enhancing the adaptability and effectiveness of the solver:
    • Memory Structure: Each entry in $M$ consists of graph attributes such as $nodes$, $edges$, $edge\_density$, $chromatic\_number$, and the corresponding $solution$.
    • Similarity Assessment: A similarity function $\sigma(x,M)$ evaluates the resemblance of the current graph attributes $x$ to those stored in $M$, prioritizing instances where attribute deviations are minimal. The similarity function is defined as:
      $\sigma(x,M)=\{(x_{j},f_{\text{v2}}(x_{j})):x_{j}\in M,\,d(x,x_{j})<\alpha\}$
      where, $d(x,x\_j)=|attributes(x)-attributes(x\_j)|$,   $\alpha$ threshold.
  • Improvement Trend Evaluation: If consecutive iterations indicate no improvement in solving capabilities, or if S1 fails to solve a problem within a set limit of five iterations, the S2 solver is invoked to ensure problem resolution.
  • Example Generation: The Graph Coloring Example Generator, $\mathcal{E}(x)$, simplifies the graph problem space by creating subgraphs. This module uses a greedy algorithm to propose potential solutions for these subgraphs, aiding the broader problem-solving strategy by focusing on simpler problem segments.
  • Feedback Generation and Validator: The Graph Coloring Validator assesses the validity of the proposed solutions and identifies errors such as coloring conflicts or uncolored vertices. Feedback denoted as $\mathcal{F}(f_{\text{S1}}(x,M_{x}))$ is provided based on these assessments, critical for the iterative adjustment of the solver’s strategy. The feedback process is formally incorporated as part of the metacognitive governance’s evaluation of $C(f_{\text{S1}}(x,M_{x}))$ against a confidence threshold $\theta_{\text{S1}}$.
This configuration ensures that SOFAI-v2 operates under a robust framework of real-time metacognitive governance, dynamically adapting to the complexity of graph coloring problems through experience-based learning and systematic feedback mechanisms. Examples of prompts, feedback types, and generated examples are in Supplementary Material.
Graph Coloring Problem Prompt New Problem to Solve:
You are given an undirected graph with 3 colors available. Your task is to assign a color to each vertex such that no two adjacent vertices share the same color.
Graph Representation:
- Number of vertices and edges: p edge 5 5.
- Edges between vertices are listed as follows: e A B
e A C
e B C
e C D
e D E
Objective:
Assign a unique color to each vertex, ensuring that no two vertices connected by an edge have the same color. Use no more than 3 distinct colors. Provide the color assignments for each vertex in the format:
(Vertex Color) Example Format:
(A 1)
(B 2)
(C 1)
Please provide the color assignment for the new problem to solve directly below:
Figure 3: LLM prompt template for $m(100,0)$ for the graph coloring problem without using episodic memory.
SOFAI-v1 vs SOFAI-v2: SOFAI-v2 introduces significant advancements over its predecessor, SOFAI-v1, by incorporating iterative feedback mechanisms and episodic memory, which enhance the problem-solving capabilities of the S1 solver. Unlike SOFAI-v1, where problems not solved by S1 are directly escalated to S2 without iterative refinement, SOFAI-v2 systematically leverages historical data and adaptive feedback to optimize solutions before resorting to S2, thereby improving efficiency and efficacy.
Episodic Memory Prompt Problem:
p edge 4 4
e X Y
e Y Z
e Z A
e X A
Correct Solution:
(X 1)
(Y 2)
(Z 1)
(A 2)
End of Example
Figure 4: Example for episodic memory, which is added at the end of LLM prompt template in Fig 8.
Performance Metrics
All solver configurations, as detailed in Table 1, were evaluated using two primary metrics: success rate and average time taken. The success rate shown in Equation 1 measures the proportion of correct solutions out of the total number of problems, while the average time shown in Equation 2 taken calculates the mean time per problem solved within the time cap. The metrics are defined as follows:
$\text{Success Rate}(\%)=\left(\frac{\text{Number of Correct Solutions}}{\text{Total Number of Problems}}\right)\times 100$ (1)
Average Time $\text{Average Time Taken}=\frac{\sum_{i=1}^{n}t_{i}}{n}$ (2)
where in Equation 2, $t_{i}$ represents the time taken to solve the $i$-th problem, and $n$ is the total number of problems solved within the time cap. Results were aggregated over three randomized trials to ensure reliability and consistency in the findings.
ID Description
S1 Mistral-7B LLM
S2 DSATUR with Backtracking
SOFAI-v1 Combines S1 and S2 with rule-based metacognition
SOFAI-v2 Combines S1 and S2 with real-time metacognition
MC-S1-I[1-5] Iterative metacognitive feedback ranging from one to five iterations to refine S1 solutions.
Table 1: Descriptions of solver configurations employed in the experiments.
Results
This section presents an analysis of solver performance by systematically controlling key parameters. Unless mentioned, the edge formation probability is fixed to $p=0.5$, allowing a focused evaluation across varying graph sizes $n$ and solvability mixes $m$ to capture a broad range of problem configurations. Tables 2 and 3 provide the success rate and average time taken for each solver configuration under these controlled conditions. The Supplementary Material includes results for additional $p$ values to supplement this focused analysis, providing a comprehensive view of solver robustness across varying edge densities. From these results, we aim to address four key research questions (RQs) that provide deeper insights into SOFAI-v2 performance and robustness.
RQ 1 What improvement in success rate does SOFAI-v2 demonstrate compared to other solvers?
Graph Size m(100,0) (in %) m(0,100) (in %) m(50,50) (in %)
S1 S2 SOFAI_v1 SOFAI-v2 S1 S2 SOFAI_v1 SOFAI-v2 S1 S2 SOFAI_v1 SOFAI-v2
5 9.41 100 100 100 75 100 100 100 45.36 100 100 100
10 0 100 100 100 60.38 100 100 100 39.39 100 100 100
15 0 80 80 80 37.50 77.08 89.58 93.75 17.65 77.45 83.33 84.31
20 0 5 5 7.29 45.65 4.35 47.83 76.09 22.58 3.06 24.73 38.95
30 0 0 0 0 54.17 0 62.50 72.92 28.89 0 33.71 38.89
40 0 0 0 0 33.33 0 47.37 47.37 17.43 0 24.55 24.77
50 0 0 0 0 3.85 0 3.85 53.85 2.11 0 2.11 27.72
Table 2: Success rates (%) across solvers and problem configurations for graph sizes from 5 to 50 with edge probability $p=0.5$. Configurations $m=(100,0)$, $m=(0,100)$, and $m=(50,50)$ represent all solvable, all unsolvable, and a balanced mix of solvable and unsolvable instances, respectively.
According to the results shown in Table 2, SOFAI-v2, SOFAI-v1, and S2 exhibit comparable performance in fully solvable problem configuration ($m=(100,0)$). In contrast, in the unsolvable scenario ($m=(0,100)$) at a graph size of 50, SOFAI-v2 achieves a success rate of 53.85%, which is significantly higher than the 3.85% achieved by both SOFAI-v1 and S1. Similarly, in the mixed scenario ($m=(50,50)$), SOFAI-v2 continues to perform well with a 27.72% success rate, far surpassing SOFAI-v1 and S1, each at 2.11%.
The higher success rates of SOFAI-v2 in challenging problem configurations likely arise from its metacognitive governance that effectively identifies and resolves constraints, particularly in unsolvable configurations. This capability allows SOFAI-v2 to adapt more dynamically to complex problem sets, enhancing its performance in environments with a mix of solvable and unsolvable problems.
Answer to RQ 1: The SOFAI-v2 model shows enhanced performance in tackling unsolvable and mixed problem configurations, surpassing other solvers as detailed in Table 2. For example, within the unsolvable scenarios set (with $m=(0,100)$ at $n=50$), SOFAI-v2 achieves a success rate of 53.85%. This is approximately a 1298% improvement over the success rates of SOFAI-v1 and Solver S1, which are around 3.85%. The calculation for the improvement is derived as follows:
$\frac{53.85-3.85}{3.85}\times 100\%\approx 1298\%$
This significant performance boost is attributed to SOFAI-v2’s metacognitive governance system, which optimizes decision-making processes in complex problem configurations.
RQ 2 How does SOFAI-v2 compare in terms of average time taken for solving problems relative to other solvers?
Graph Size m(100,0) (in seconds) m(0,100) (in seconds) m(50,50) (in seconds)
S1 S2 SOFAI_v1 SOFAI-v2 S1 S2 SOFAI_v1 SOFAI-v2 S1 S2 SOFAI_v1 SOFAI-v2
5 3.23 0.001 3.23 0.92 3.13 0.001 3.13 0.77 3.23 0.001 3.23 0.92
10 4.56 0.19 4.69 2.21 4.08 0.17 4.14 1.07 5.28 0.22 5.48 3.74
15 5.61 70.85 55.46 53.23 6.20 72.02 39.89 22.33 5.12 67.08 70.01 85.37
20 6.79 199.27 154.11 120.86 7.52 198.70 111.00 46.96 6.68 196.22 193.15 195.25
30 11.53 200.00 137.80 117.10 10.87 200.00 82.30 52.10 9.92 200.00 192.85 193.67
40 9.69 200.00 153.41 144.54 9.46 200.00 109.71 101.18 10.30 200.00 194.47 193.67
50 41.98 200.00 200.00 139.17 62.76 200.00 88.93 88.93 21.83 200.00 200.00 194.47
Table 3: Average time taken (in seconds) for different solvers across graph sizes and problem configurations, highlighting operational efficiency in solving or recognizing unsolvable problems.
According to the results in Table 3, SOFAI-v2 demonstrates distinct time efficiency advantages across various problem configurations and graph sizes. In fully solvable problems ($m=(100,0)$), the average solving time for SOFAI-v2 remains consistent and competitive with other solvers. However, in more complex problem configurations, such as all unsolvable problems ($m=(0,100)$) at $n=50$, SOFAI-v2 significantly reduces the average solving time to 88.93 seconds, a 55% decrease compared to S2, requiring up to 200 seconds. Similarly, in the mixed scenario ($m=(50,50)$) at the same graph size, SOFAI-v2 further improves efficiency, reducing the solving time to 194.47 seconds, demonstrating a nearly 3% improvement over SOFAI_v1 (200 seconds) and S2 (200 seconds).
The enhanced time efficiency of SOFAI-v2 in these challenging configurations likely results from its metacognitive governance, which optimizes computational strategies to handle complex and mixed problem configurations more efficiently. This capability enables SOFAI-v2 to solve problems faster and manage its computational resources more effectively, particularly in problem configurations where solvability varies widely.
Answer to RQ 2: In terms of time efficiency, SOFAI-v2 significantly outperforms other solvers, notably in problem configurations involving unsolvable and mixed problems. Specifically, at a graph size of 50 within the unsolvable configuration ($m=(0,100)$), the average time taken by SOFAI-v2 to attempt solving these problems is 88.93 seconds, which is approximately 55% less than the 200 seconds typically required by Solver S2. The percentage reduction is calculated as follows:
$\frac{200-88.93}{200}\times 100\%\approx 55\%$
This enhanced performance is attributed to the metacognitive governance within SOFAI-v2, which accelerates the problem-solving process and increases accuracy in complex problem configurations.
RQ 3 Does the iterative calling mechanism in SOFAI-v2 enhance the performance of its S1 solver through metacognitive governance?
(a) All Solvable ($m=(100,0)$)
(b) All Unsolvable ($m=(0,100)$)
(c) Balanced Mix ($m=(50,50)$)
Figure 5: Success rates of SOFAI-v2’s S1 with iterative metacognitive feedback (MC-S1-I1 to MC-S1-I5) across different graph sizes and problem configurations, showing the impact of iterations on solver performance.
(a) All Solvable ($m=(100,0)$)
(b) All Unsolvable ($m=(0,100)$)
(c) Balanced Mix ($m=(50,50)$)
Figure 6: Success rates across solvers as a function of edge probability in different problem configurations, highlighting the adaptability of SOFAI-v2 compared to traditional solvers.
(a) All Solvable ($m=(100,0)$)
(b) All Unsolvable ($m=(0,100)$)
(c) Balanced Mix ($m=(50,50)$)
Figure 7: Time efficiency across solvers as a function of edge probability in different problem configurations, showcasing SOFAI-v2’s performance in reducing average time taken.
The iterative feedback mechanism within SOFAI-v2’s S1 solver shows a clear progression in performance across different graph sizes and problem configurations, as detailed in Figure 5. Each iteration, from MC-S1-I1 through MC-S1-I5, demonstrates a quantitative improvement in success rates, indicative of the solver’s enhanced problem-solving capabilities due to metacognitive governance.
In the fully solvable problems ($m=(100,0)$), the success rates of the initial iteration MC-S1-I1 start at approximately 20% and quickly improve to nearly 80% by MC-S1-I5 for $n=5$. This rapid improvement highlights the effectiveness of iterative feedback in optimizing the solver’s approach to solvable problems.
For the entirely unsolvable problems ($m=(0,100)$), success rates at $n=10$ begin at about 60% during MC-S1-I1 and rise to over 85% by MC-S1-I5. This increase of approximately 40% in success rate across iterations underscores the solver’s growing capability to efficiently recognize and handle unsolvable problem configuration without unnecessarily invoking higher-cost computational resources.
In the balanced mix problem configuration ($m=(50,50)$), iterative feedback leads to substantial improvement across graph sizes. Initially, success rates at MC-S1-I1 are around 40% for $n=10$ and increase to about 60% by MC-S1-I5. This demonstrates a consistent improvement of around 50% in discerning between solvable and unsolvable problems as iterations progress, illustrating the benefits of metacognitive governance in mixed problem configurations.
Answer to RQ 3: The iterative calling mechanism of SOFAI-v2 substantially enhances the performance of its S1 solver in various problem configurations. For instance, in the configuration where $m=(100,0)$, success rates escalate from 20% in the MC-S1-I1 to nearly 80% by the MC-S1-I5, marking an approximate 300% increase. This improvement is calculated as follows:
$\frac{80-20}{20}\times 100\%=300\%$
Similarly, in the configuration $m=(0,100)$, the success rates rise by 40%, from 60% in MC-S1-I1 to 85% in MC-S1-I5:
$\frac{85-60}{60}\times 100\%\approx 40\%$
In the balanced scenario ($m=(50,50)$), there is a 50% increase in success rates, from 40% in MC-S1-I1 to 60% in MC-S1-I5, demonstrating the solver’s adaptability:
$\frac{60-40}{40}\times 100\%=50\%$
The metacognitive governance within SOFAI-v2 facilitates these significant improvements across different configurations, enhancing the solver’s efficiency and adaptability.
RQ4 How does edge probability ($p$) influence the solvability and time efficiency of graph coloring problems across different solvers?
The impact of edge probability ($p$) on solvability and time efficiency across solvers was systematically evaluated through success rates and solving times as $p$ ranges from 0.1 to 0.9. As $p$ increases, all solvers generally exhibit a decreasing trend in success rates until reaching a plateau, as detailed in Figure 6. This trend indicates the increasing complexity of the graph coloring problems with more connections between nodes. Regarding time efficiency, there is a noticeable increase in average solving time, peaking before stabilizing, as shown in Figure 7.
SOFAI-v2 consistently outperforms S1, S2, and SOFAI-v1 in maintaining higher success rates across all edge probabilities, demonstrating enhanced resilience against increasing problem complexity. Additionally, SOFAI-v2 consistently requires less time to solve problems across all configurations, highlighting its superior time efficiency.
In the All Solvable ($m=(100,0)$) configuration, SOFAI-v2 exhibits comparable success rates as the probability $p$ increases, similar to other solvers. In the All Unsolvable ($m=(0,100)$) scenario, SOFAI-v2 achieves peak success rates at an edge probability of $p=0.1$, which are 20-40% higher than other solvers (SOFAI-v2 = 94%, SOFAI-v1 = 74%, Solver S1 = 68%, Solver S2 = 51%). This advantage is maintained with an approximate 10-20% lead as $p$ increases, as illustrated in Figure 6. Additionally, in this configuration, SOFAI-v2 solves problems 60-80% faster than its counterparts at mid-range probabilities ($p=0.1$ to $0.5$). For instance, at $p=0.1$, the solving times are as follows: SOFAI-v2 = 20 seconds, SOFAI-v1 = 55 seconds, and Solver S2 = 112 seconds, as depicted in Figure 7. In the Balanced Mix ($m=(50,50)$) scenario, the success rate for SOFAI-v2 peaks at approximately 72% at $p=0.1$, which is about 15-25% higher than the rates observed for SOFAI-v1, Solver S1, and Solver S2. The time efficiency in this scenario also notably surpasses other solvers, with reductions in solving times of 40-50%. Specifically, at $p=0.1$, the times are: SOFAI-v2 = 70 seconds, SOFAI-v1 = 100 seconds, and Solver S2 = 108 seconds.
Answer to RQ 4: Edge probability ($p$) substantially influences solvability and time efficiency, with SOFAI-v2 showing superior adaptability and operational efficiency. In problem configurations with increasing complexity, SOFAI-v2 sustains higher success rates and achieves significant time efficiency compared to other solvers, affirming its effective metacognitive governance.
Discussion
This study presented SOFAI-v2, a neurosymbolic architecture that integrates a fast, experience-based S1 powered by an LLM with a slow, deliberative S2 utilizing the DSATUR graph coloring algorithm governed by real-time metacognitive governance. Our experiments evaluated the performance of SOFAI-v2 on graph coloring problems with sizes ranging from 5 to 50, generated using the Erdős–Rényi model with an edge probability of $p=0.5$. We assessed the solvers across three problem configurations: all solvable ($m=(100,0)$), all unsolvable ($m=(0,100)$), and a balanced mix ($m=(50,50)$). The discussion addresses the research questions posed.
Integrating metacognitive governance and episodic memory in SOFAI-v2 demonstrates a notable enhancement in solving graph coloring problems, significantly improving success rates and reducing computational times across various graph sizes and problem complexities. Success rates in complex scenarios like the all-unsolvable configuration reveal SOFAI-v2’s superior adaptability, achieving up to a 53.85% success rate for larger graphs, a substantial improvement over SOFAI-v1 and traditional symbolic solvers. Additionally, iterative metacognitive feedback in SOFAI-v2’s System 1 has proven effective in refining problem-solving strategies, increasing efficiency, and enhancing the solver’s ability to differentiate between solvable and unsolvable problems. These capabilities underscore the value of combining neural and symbolic approaches to overcome limitations in traditional CSP solvers, demonstrating SOFAI-v2’s potential in dynamic and demanding CSP environments.
Implications and Comparisons with Prior Work
This study broadens the scope of neurosymbolic systems by applying them to graph coloring problems. It demonstrates that integrating LLMs with traditional symbolic solvers, guided by metacognitive governance, significantly enhances solutions’ accuracy and efficiency. Unlike previous applications that primarily leveraged LLMs for sequential decision-making tasks (Valmeekam et al. 2022; Stechly, Valmeekam, and Kambhampati 2024), SOFAI-v2 directly addresses their constraint adherence limitations through iterative feedback and strategic use of symbolic reasoning. This approach aligns with cognitive decision-making theories and shows a better success rate and time efficiency, illustrating the practical benefits of hybrid architectures in complex constraint satisfaction problems.
Conclusion
We introduced SOFAI-v2, a neurosymbolic architecture that integrates a fast, episodic memory-based LLM (S1) with a slow, deliberative symbolic solver (S2), enhanced by real-time metacognitive governance. Our empirical evaluations across various graph coloring problems confirm that SOFAI-v2 outperforms traditional symbolic solvers and its predecessor, SOFAI-v1, in both success rates and time efficiency, especially within complex problem configurations.
SOFAI-v2 employs iterative metacognitive feedback and episodic memory-based iterations to enable S1 to refine its response. This effectively overcomes the limitations of LLMs in constraint adherance of CSPs. By leveraging episodic memory and adaptive feedback, SOFAI-v2 significantly enhances the accuracy and efficiency of solving CSPs, demonstrating the potential of cognitive-inspired AI architectures in managing complex challenges. Future work will scale SOFAI-v2 to tackle larger and more varied problem sets, explore different graph models, and apply the architecture to other types of CSPs and sequential decision-making tasks. Enhancing the metacognitive components with advanced learning algorithms and incorporating alternative neural models are anticipated to boost its adaptability and efficiency further, contributing to developing robust, scalable solutions for intricate AI applications.
Appendix A Prompt, Example, and Feedback Examples
This section discusses prompt type, feedback types, and generated examples. Feedback and generated example types are examples of the implementation of SOFAI-v2 iterative feedback mechanism with example generation.
Graph Coloring Problem Prompt New Problem to Solve:
You are given an undirected graph with 2 colors available. Your task is to assign a color to each vertex such that no two adjacent vertices share the same color.
Graph Representation:
- Number of vertices and edges: p edge 5 5.
- Edges between vertices are listed as follows: e A B
e A C
e B C
e C D
e D E
Objective:
Assign a unique color to each vertex, ensuring that no two vertices connected by an edge have the same color. Use no more than 2 distinct colors. Provide the color assignments for each vertex in the format:
(Vertex Color) Example Format:
(A 1)
(B 2)
(C 1)
Please provide the color assignment for the new problem to solve, or respond with ”NOT SOLVABLE” if it cannot be solved.
Figure 8: LLM prompt template for $m(0,100)$ and $m(50,50)$ for the graph coloring problem without using episodic memory.
Graph Coloring Problem Prompt with Episodic Memory Example New Problem to Solve:
You are given an undirected graph with 2 colors available. Your task is to assign a color to each vertex such that no two adjacent vertices share the same color.
Graph Representation:
- Number of vertices and edges: p edge 5 5.
- Edges between vertices are listed as follows: e A B
e A C
e B C
e C D
e D E
Objective:
Assign a unique color to each vertex, ensuring that no two vertices connected by an edge have the same color. Use no more than 2 distinct colors. Provide the color assignments for each vertex in the format:
(Vertex Color) Episodic Memory Example:
Graph:
p edge 3 3
Edges:
e A B
e B C
e C A
Solution:
(A 1)
(B 2)
(C 3)
This example is provided to demonstrate a previously successful coloring strategy for a similar subgraph, which may be helpful in solving the current problem. Please provide the color assignment for the new problem to solve directly below, or respond with ”NOT SOLVABLE” if it cannot be solved.
Figure 9: LLM prompt template for $m(0,100)$ and $m(50,50)$ for the graph coloring problem, incorporating an episodic memory example to aid problem-solving.
Graph Coloring Feedback - Exceeding Limit An attempt was made to solve the graph coloring using three colors, exceeding the maximum allowed. Over-Colored Submission:
(A 1)
(B 2)
(C 3)
(D 1)
(E 2)
Feedback Provided:
Error: Only 2 colors are allowed. 3 colors were used.
Figure 10: Examples of feedback provided to coloring assignment, exceeding the number of colors limit.
Graph Coloring Feedback – Incorrect Coloring Given an undirected graph with 2 colors available, the vertices were incorrectly colored as follows: Graph Representation:
- Number of vertices and edges: p edge 5 5.
- Edges between vertices are listed as follows: e A B
e A C
e B C
e C D
e D E
Incorrect Coloring Submitted:
(A 1)
(B 1)
(C 2)
(D 2)
(E 1)
Feedback Provided:
Error: Vertices A and B are adjacent but have the same color.
Error: Vertices B and C are adjacent but have the same color.
Figure 11: Examples of feedback provided to incorrect coloring assignment.
Graph Coloring Generated Example To generate an example, a subgraph is extracted from the problem graph to demonstrate a valid coloring with fewer vertices. Subgraph Representation:
- Subgraph vertices and edges: p edge 3 3.
- Edges between subgraph vertices are listed as follows: e A B
e A C
e B C
Correct Coloring of Subgraph:
(A 1)
(B 2)
(C 3)
This example shows how to correctly assign colors in a smaller scope of the main problem. Scaling this approach for larger graphs while maintaining the constraints is crucial.
Figure 12: Example of a correctly colored subgraph is generated as part of the feedback to illustrate the correct coloring assignment for less number of nodes.
Appendix B Results - Success Rate
This section explores the success rates of different solvers in solving graph coloring problems across different problem configurations ($m$). Figure 13, Figure 14 and Figure 15 demonstrate how each solver—System 1, System 2, SOFAI-v1 and SOFAI-v2 performs across varying edge probabilities ($p$). These graphs show the robustness of SOFAI-v2 in terms of success rate compared to other solvers.
Figure 13: Success Rate (SR) of different solvers vs Graph size ($n$) across edge probabilities ($p$) for problem configuration ($m=(50,50)$)
Figure 14: Success Rate (SR) of different solvers vs Graph size ($n$) across edge probabilities ($p$) for problem configuration ($m=(0,100)$)
Figure 15: Success Rate (SR) of different solvers vs Graph size ($n$) across edge probabilities ($p$) for problem configuration ($m=(100,0)$)
Appendix C Results - Average Time Take
This section explores the time taken by different solvers to solve graph coloring problems across different problem configurations ($m$). Figure 16, Figure 17, and Figure 18 demonstrate how each solver—System 1, System 2, SOFAI-v1, and SOFAI-v2—performs across varying edge probabilities ($p$). These graphs show the time efficiency of SOFAI-v2 compared to System 2 and SOFAI-v1.
Figure 16: Average time of different solvers vs Graph size ($n$) across edge probabilities ($p$) for problem configuration ($m=(50,50)$)
Figure 17: Average time of different solvers vs Graph size ($n$) across edge probabilities ($p$) for problem configuration ($m=(0,100)$)
Figure 18: Average time of different solvers vs Graph size ($n$) across edge probabilities ($p$) for problem configuration ($m=(100,0)$)

### v2

@@
 Existing symbolic solvers are often slow, and prior research has shown that Large Language Models (LLMs) alone struggle with CSPs because of their complexity.
-To bridge this gap, we build upon the existing SOFAI architecture (or SOFAI-v1), which adapts Daniel Kahneman’s “Thinking, Fast and Slow” cognitive model to AI.
-Our enhanced architecture, SOFAI-v2, integrates refined metacognitive governance mechanisms to improve adaptability across complex domains, specifically tailored here for solving CSPs like graph coloring.
-SOFAI-v2 combines a fast System 1 (S1) based on LLMs with a deliberative System 2 (S2) governed by a metacognition module.
-S1’s initial solutions, often limited by non-adherence to constraints, are enhanced through metacognitive governance, which provides targeted feedback and examples to adapt S1 to CSP requirements.
-If S1 fails to solve the problem, metacognition strategically invokes S2, ensuring accurate and reliable solutions.
-With empirical results, we show that SOFAI-v2 for graph coloring problems achieves an 16.98% increased success rate and is 32.42% faster than symbolic solvers.
-Introduction
-Constraint Satisfaction Problems are a core challenge in artificial intelligence (AI) due to their demand for correctness adhering to strict constraints in static (Kumar 1992) or uncertain, dynamic environments (Verfaillie and Jussien 2005).
-These problems are prominent in applications such as scheduling, boolean satisfiability, resource allocation, temporal reasoning, and planning (Welsh and Powell 1967; Chaitin 1982).
-Traditional approaches to solve CSPs typically rely on constraint propagation or search (Haralick and Elliott 1980).
-The symbolic approaches are known for their accuracy but often struggle with scalability as the complexity of CSPs increases, hindered by high computational overhead and slow processing times (Dechter 2003; Rossi, Van Beek, and Walsh 2006).
-Recent advancements in LLMs have shown promise in rapidly processing complex information (Ruoss et al.
-2024), but they are fundamentally limited in solving sequential decision-making tasks like planning (Valmeekam et al.
-2022), and reasoning (Stechly, Valmeekam, and Kambhampati 2024).
-Due to their probabilistic, retrieval-based nature, LLMs approximate solutions rather than generating definitive answers, often leading to partial or inconsistent adherence to constraints (Jiang et al.
-2023).
-Even with state-of-the-art prompting strategies and self-verification techniques, LLMs fall short in handling the strict requirements of CSPs, making their outputs unreliable for applications demanding high precision.
-The disparity between the precision of traditional symbolic methods and the adaptability and speed of LLMs underscores the need for a new paradigm in tackling complex tasks like graph coloring (Jensen and Toft 2011), a canonical CSP.
-Drawing inspiration from Daniel Kahneman’s cognitive theory of “Thinking, Fast and Slow” (Kahneman 2011) and its adaptation to AI as SOFAI-v1 (Booch et al.
-2021), we introduce SOFAI-v2 which extends the earlier architecture by refining metacognitive mechanisms that oversee a fast, experience-based S1, powered by an LLM, and a slow, deliberative S2 using the DSATUR algorithm for graph coloring.
+To bridge this gap, we build upon the existing SOFAI architecture (SOFAI-v1), which adapts Daniel Kahneman’s “Thinking, Fast and Slow” cognitive model to AI.
+Our enhanced architecture, SOFAI-v2, integrates refined metacognitive governance mechanisms to improve adaptability across complex domains, specifically tailored here for solving the graph coloring problem, a specific type of CSP.
+SOFAI-v2 combines a fast System 1 (S1), leveraging LLMs, with a deliberative System 2 (S2), governed by a metacognition module.
+S1’s initial solutions, often limited by constraint adherence issues, are improved through targeted feedback and examples from metacognition, aligning S1 more closely with CSP requirements.
+If S1 fails to resolve the problem, metacognition strategically invokes S2, ensuring accurate and reliable solutions.
+Our empirical results demonstrate that SOFAI-v2 achieves a 10.5% higher success rate and is up to 30% faster than a traditional symbolic solver in solving graph coloring problems.
+vedant@email.sc.edu vishalp@mailbox.sc.edu biplav.s@sc.edu francesca.rossi2@ibm.com
+1University of South Carolina, Columbia, SC, USA.
+2IBM Research, NY, USA.
+1 Introduction
+Constraint Satisfaction Problems are a core challenge in artificial intelligence (AI) due to their demand for correctness adhering to strict constraints in static (Kumar, 1992) or uncertain, dynamic environments (Verfaillie & Jussien, 2005).
+These problems are prominent in applications such as scheduling, boolean satisfiability, resource allocation, temporal reasoning, and planning (Welsh & Powell, 1967; Chaitin, 1982).
+Traditional approaches to solve CSPs typically rely on constraint propagation or search (Haralick & Elliott, 1980).
+The symbolic approaches are known for their accuracy but often struggle with scalability as the complexity of CSPs increases, hindered by high computational overhead and slow processing times (Dechter, 2003; Rossi et al., 2006).
+Recent advancements in LLMs have shown promise in rapidly processing complex information (Ruoss et al., 2024), but they are fundamentally limited in solving sequential decision-making tasks like planning (Valmeekam et al., 2022), and reasoning (Stechly et al., 2024).
+Due to their probabilistic, retrieval-based nature, LLMs approximate solutions rather than generating definitive answers, often leading to partial or inconsistent adherence to constraints (Jiang et al., 2023).
+Even with state-of-the-art prompting strategies and self-verification techniques, LLMs fall short in handling strict requirements of CSPs, making their outputs unreliable for applications demanding high precision.
+We study feasibility under fixed resource limits in the graph coloring decision problem, a core constraint satisfaction setting that supports clear verification and iterative guidance.
+The disparity between the precision of traditional symbolic methods and the adaptability and speed of LLMs underscores the need for a new paradigm in tackling complex tasks like graph coloring (Jensen & Toft, 2011), a canonical CSP.
+Drawing inspiration from Daniel Kahneman’s cognitive theory of “Thinking, Fast and Slow” (Kahneman, 2011) and its adaptation to AI as SOFAI-v1 (Fabiano et al., 2025; Booch et al., 2021), we introduce SOFAI-v2 which extends the earlier architecture by refining metacognitive mechanisms that oversee a fast, experience-based S1, powered by an LLM, and a slow, deliberative S2 using the DSATUR algorithm for graph coloring.
 This governance mechanism continuously monitors and enhances S1’s outputs, offering targeted feedback and constraint adherence examples; when S1’s solutions fall short, S2 is strategically invoked to guarantee accuracy.
 In this setup, S1 swiftly produces initial solutions, while S2 provides a reliable fallback for achieving precise constraint satisfaction.
-Our contributions are as follows:
-  • We introduce SOFAI-v2, a neurosymbolic fast and slow architecture designed to enhance the performance of CSP solvers on graph coloring problems through adaptive metacognitive governance.
-  • We develop a comprehensive benchmark suite of graph coloring problems featuring variations in graph size, edge probability, and a mix of solvable and unsolvable instances to test solver adaptability.
-  • We conduct extensive empirical evaluations of various solver configurations for graph coloring, demonstrating the advantages of SOFAI-v2 in terms of success rate, time efficiency, and adaptability.
+This work studies problem solving with a fast S1, a deliberate S2, and an MC that monitors and guides iterations, with memory for reuse.
+Beyond the S1/S2 framing, SOFAI-v2 is positioned as a neurosymbolic cognitive architecture: S1 (neural) proposes, symbolic constraints verify, and MC enacts control.
+This aligns with computational accounts of automatic versus controlled processing and supervisory control (e.g., contention scheduling and supervisory attention) (Norman & Shallice, 1986; Cooper & Shallice, 2000), providing a bridge from cognitive theory to an AI system.
+The implemented system targets the graph coloring decision task and evaluates success rate, time, and iteration behavior. Our contributions are as follows:
+  • Introduced SOFAI-v2, a neurosymbolic fast and slow architecture to enhance the performance of CSP solvers on graph coloring problems through adaptive metacognitive governance.
+  • Created a comprehensive benchmark of graph coloring problems, featuring variations in graph size, edge probability, and solvable and unsolvable instances to test adaptability.
+  • Conducted extensive empirical evaluations of various solver configurations for graph coloring, demonstrating advantages of SOFAI-v2 in terms of success rate, time efficiency.
 The rest of the paper is organized as follows: we first provide the necessary background and discuss prior work.
@@
 Finally, we conclude with a summary of our findings.
-Background
-Constraint Satisfaction Problem
-CSP is defined as a tuple $(X,D,C)$, where:
-  • $X=\{x_{1},x_{2},\dots,x_{n}\}$ is a finite set of variables.
-  • $D=\{D_{1},D_{2},\dots,D_{n}\}$ is a set of domains, with each $D_{i}$ representing the possible values for variable $x_{i}$.
-  • $C=\{c_{1},c_{2},\dots,c_{m}\}$ is a set of constraints, with each constraint $c_{j}$ defining a subset of allowed assignments for a set of variables involved in the constraint.
-A solution to a CSP is an assignment of values to all variables in $X$ such that every constraint in $C$ is satisfied.
-CSPs are used extensively in scheduling, resource allocation, and combinatorial optimization, where satisfying multiple constraints simultaneously is crucial.
-Graph Coloring Problem
-The graph coloring problem is a specific type of CSP, where the goal is to assign a color to each node of a graph such that no two adjacent nodes share the same color.
-Formally, it can be defined as a tuple $(G,C,k)$, where:
-  • $G=(V,E)$ is an undirected graph, with $V=\{v_{1},v_{2},\dots,v_{n}\}$ representing the set of vertices (or nodes) and $E=\{(v_{i},v_{j})|v_{i},v_{j}\in V\}$ representing the set of edges.
-  • $C=\{c_{1},c_{2},\dots,c_{k}\}$ is a set of $k$ colors.
-  • The objective is to find a function $f:V\rightarrow C$ such that for any edge $(v_{i},v_{j})\in E$, $f(v_{i})\neq f(v_{j})$.
-We use the DIMACS representation (Johnson and Trick 1996) to describe the graph coloring problem, and its corresponding visualization is shown in Figure 1.
-1 c Example of a graph coloring problem in DIMACS format
-2 p edge 5 5
-3 c edges
-4 e A B
-5 e A C
-6 e B C
-7 e C D
-8 e D E
-(a) DIMACS Format of Graph Coloring Problem
-(b) Visualization of the Graph
-Figure 1: Graph Coloring Problem in DIMACS Format and Its Visualization
-Degree of Saturation Algorithm with Backtracking
-The Degree of Saturation, or DSATUR algorithm is a heuristic for approaching optimal coloring by assigning colors to vertices based on their saturation degree—the number of distinct colors already assigned to adjacent vertices (Brélaz 1979).
-While the algorithm by itself is not guaranteed to find an optimal solution, the addition of backtracking allows it to systematically explore alternative color assignments, ensuring optimal coloring is achieved.
-Given the graph coloring problem $(G,C,k)$, where $G=(V,E)$ is an undirected graph, $C=\{c_{1},c_{2},\dots,c_{k}\}$ is a set of colors, and $f:V\rightarrow C$ is the coloring function, the algorithm proceeds as follows:
-  1. 1.
-    Initialization: Set $f(v)=\emptyset$ for all $v\in V$, marking all vertices as uncolored.
-  2. 2.
-    Saturation Degree Calculation: For each uncolored vertex $v\in V$, calculate its saturation degree, $d_{sat}(v)$, defined as the number of distinct colors assigned to its adjacent vertices: $\{f(u)\mid(v,u)\in E,f(u)\neq\emptyset\}$.
-  3. 3.
-    Vertex Selection: Choose the uncolored vertex $v$ with the highest saturation degree $d_{sat}(v)$.
-In the case of a tie, select the vertex with the highest degree $d(v)$, which is the number of edges incident to $v$.
-  4. 4.
-    Color Assignment: Assign the smallest feasible color $c\in C$ to $v$, ensuring that $f(v)\neq f(u)$ for all adjacent vertices $(v,u)\in E$.
-  5. 5.
-    Backtracking (to Ensure Optimality): If no valid color can be assigned without violating constraints, initiate backtracking.
-Backtracking systematically undoes recent assignments and explores alternate color choices.
-This recursive search allows the algorithm to revisit decisions, thereby exploring all possible color assignments and guaranteeing that an optimal solution—one using the minimal number of colors—is achieved.
-  6. 6.
-    Repeat: Continue steps 2-5 until all vertices are colored or no valid coloring exists.
-The algorithm finds a minimal coloring assignment by prioritizing vertices with the highest constraints and employing backtracking to explore alternatives.
-Without backtracking, the heuristic alone may fall short of optimality; with backtracking, however, the algorithm can reach an optimal solution.
-Related Works
-Figure 2: SOFAI-v2 architecture with real-time metacognitive governance
+2 Background
+2.1 Graph Coloring Decision Problem
+The graph coloring decision problem is a specific type of constraint satisfaction problem (CSP), where the objective is to determine whether a given graph can be colored using at most $k$ colors such that no two adjacent nodes share the same color.
+Formally, the problem is defined as follows:
+  • An undirected, unweighted graph $G=(V,E)$
+  • A positive integer $k$ representing the maximum number of colors
+  • Determine whether there exists a function $f:V\rightarrow\{1,2,\dots,k\}$ such that for every edge $(u,v)\in E$, $f(u)\neq f(v)$.
+Unlike the classical graph coloring problem, which aims to minimize the number of colors used (an optimization problem), the decision version focuses solely on satisfiability: whether a valid coloring exists for a given $k$.
+We use the DIMACS representation (Johnson & Trick, 1996) to describe the graph coloring problem.
+The decision problem asks a clear yes/no question for a fixed $k$.
+This matches many practical settings where a fixed resource limit is given and the goal is feasibility, e.g., schedules with a fixed number of time slots, register allocation with a fixed number of registers, or frequency assignment with a fixed set of bands.
+It provides MC with a simple accept/reject signal and makes it straightforward to generate targeted feedback when constraints fail, which we interpret as a cue for supervisory control in the architecture (Norman & Shallice, 1986; Cooper & Shallice, 2000).
+2.2 Degree of Saturation Algorithm with Backtracking
+The Degree of Saturation (DSATUR) algorithm assigns colors to graph vertices based on their saturation degree, prioritizing vertices that are the most constrained—those having adjacent vertices with the greatest variety of colors already assigned (Brélaz, 1979).
+DSATUR remains competitive for graph coloring due to its heuristic efficiency, especially for sparse and mid-sized graphs, and its direct operation on graph structures without requiring transformations typical of SAT solvers.
+Recent studies (Yekezare et al., 2024; Schidler & Szeider, 2023) highlight its effectiveness, making it a suitable choice for our symbolic solver.
+However, the traditional DSATUR algorithm targets the optimization version of the graph coloring problem, minimizing the number of colors.
+For our work, we adapt DSATUR into a satisfiability solver by constraining it with a fixed maximum number of colors.
+A detailed description of the modified algorithm is provided in Supplementary Material Section 7.
+Unlike other solvers, DSATUR can be straightforwardly modified for the decision version of the problem, making it a practical and adaptable choice for our symbolic backend.
+In SOFAI-v2, DSATUR with backtracking plays the role of slow, controlled processing that is selectively invoked by MC under conflict or low confidence, consistent with supervisory control accounts in cognitive systems (Norman & Shallice, 1986; Cooper & Shallice, 2000).
+3 Related Works
+Figure 1: SOFAI-v2 architecture with real-time metacognitive governance (supervisory control over fast/automatic and slow/controlled processing)
 This section reviews the evolution of solvers for graph coloring, the application of LLMs in sequential decision-making tasks, and the emergence of neurosymbolic approaches for enhancing decision-making processes.
-Solvers for CSP
+We also situate SOFAI-v2 within cognitive control theories that distinguish routine (automatic) processing from controlled, supervised processing (Norman & Shallice, 1986; Cooper & Shallice, 2000).
+3.1 Solvers for Graph Coloring
 Traditional symbolic solvers have been extensively developed for this problem, leveraging algorithms that systematically explore the search space.
-Early works like Brélaz (1979) introduced efficient heuristics such as the DSATUR algorithm, significantly improving practical performance.
-Chaitin (1982) applied graph coloring to compiler optimization, demonstrating the versatility of symbolic methods.
-Golumbic (2004) provided a comprehensive treatment of graph algorithms, including coloring techniques.
-Despite their accuracy, symbolic approaches often struggle with scalability as CSP complexity increases, hindered by high computational overhead and slow processing times (Dechter 2003).
-Techniques like constraint propagation (Mackworth 1977) and backtracking algorithms (Bitner and Reingold 1975) have been employed to enhance efficiency but are limited by the exponential growth of the search space.
+Early works like (Brélaz, 1979) introduced efficient heuristics such as the DSATUR algorithm, significantly improving practical performance.
+(Chaitin, 1982) applied graph coloring to compiler optimization, demonstrating the versatility of symbolic methods.
+(Golumbic, 2004) provided a comprehensive treatment of graph algorithms, including coloring techniques.
+Despite their accuracy, symbolic approaches often struggle with scalability as CSP complexity increases, hindered by high computational overhead and slow processing times (Dechter, 2003).
+Techniques like constraint propagation (Mackworth, 1977) and backtracking algorithms (Bitner & Reingold, 1975) have been employed to enhance efficiency but are limited by the exponential growth of the search space.
+DSATUR selects at each step the uncolored vertex with the highest saturation degree (the number of distinct colors among its colored neighbors), breaking ties by degree (Brélaz, 1979).
+This ordering aggressively focuses search on the most constrained vertices, yielding strong practical performance.
+With appropriate data structures, per-step selection can be done in near $O(logn)$ update time, but the overall search with backtracking remains exponential in the worst case due to the NP-completeness of graph coloring.
+In our setting, we couple DSATUR with backtracking under a fixed color budget $k$: the procedure attempts to extend a partial assignment using at most $k$ colors; failure to extend yields a certified UNSAT outcome, while success produces a constructive $k$-coloring witness.
 To address these limitations, neural solvers have been explored, utilizing machine learning techniques to approximate solutions for graph coloring.
-For instance, Li, Chen, and Koltun (2018) proposed using graph neural networks to tackle combinatorial optimization problems.
-Similarly, Khalil et al.
-(2017) introduced a framework that learns heuristics directly from data, improving scalability.
+For instance, (Li et al., 2018) proposed using graph neural networks to tackle combinatorial optimization problems.
+Similarly, (Khalil et al., 2017) introduced a framework that learns heuristics directly from data, improving scalability.
 However, neural approaches often face challenges in strictly adhering to the constraints inherent in CSPs.
-LLMs for Reasoning Tasks
+In contrast, our SOFAI setup directly targets the decision form of graph coloring with an explicit $k$, integrating DSATUR+backtracking as S2 and using S1/SOFAI components to propose assignments that are subsequently verified.
+Many end-to-end neural solvers are not readily adaptable to this satisfiability variant, which limits their comparability to our setting; we therefore retain them here primarily for historical context and contrast, while focusing our technical discussion on DSATUR and its role in our pipeline.
+3.2 LLMs for Reasoning Tasks
 Recent developments in LLMs have opened new avenues for addressing sequential decision-making tasks, including CSPs.
-Various prompt engineering techniques have been proposed to enhance the reasoning capabilities of LLMs (Brown 2020; Wei et al.
-2022; Kojima et al.
-2022; Nye et al.
-2021; Wang et al.
-2022).
-Han et al.
-(2023) demonstrated that LLMs can arrive at correct answers for graph-based problems through iterative prompting.
-However, the reasoning capabilities of LLMs have been found to be fundamentally approximate retrieval in nature (Kambhampati 2024), limiting their effectiveness in precise reasoning tasks.
-This limitation is evidenced by studies showing LLMs’ inability to solve planning problems (Valmeekam et al.
-2022) and graph coloring problems (Stechly, Valmeekam, and Kambhampati 2024), where strict adherence to constraints is required (Jiang et al.
-2023).
-Neurosymbolic Approaches for Reasoning Tasks
-Neurosymbolic approaches have gained considerable attention in addressing complex sequential decision-making tasks such as automated planning (Fabiano et al.
-2023), constrained grid navigation (Ganapini et al.
-2022), and puzzle solving (Lin et al.
-2024).
+Various prompt engineering techniques have been proposed to enhance the reasoning capabilities of LLMs (Brown, 2020; Wei et al., 2022; Kojima et al., 2022; Nye et al., 2021; Wang et al., 2022).
+(Han et al., 2023) demonstrated that LLMs can arrive at correct answers for graph-based problems through iterative prompting.
+However, the reasoning capabilities of LLMs have been found to be fundamentally approximate retrieval in nature Kambhampati (2024), limiting their effectiveness in precise reasoning tasks.
+This limitation is evidenced by studies showing LLMs’ inability to solve planning problems (Valmeekam et al., 2022) and graph coloring problems (Stechly et al., 2024), where strict adherence to constraints is required (Jiang et al., 2023).
+In cognitive-systems terms, LLM-only routines resemble contention among habitual schemas without sufficient supervisory control (Norman & Shallice, 1986).
+3.3 Neurosymbolic Approaches for Reasoning Tasks
+Neurosymbolic approaches have gained considerable attention in addressing complex sequential decision-making tasks such as automated planning (Fabiano et al., 2023), constrained grid navigation (Ganapini et al., 2022), and puzzle solving (Lin et al., 2024).
 These studies demonstrate that integrating neural methods, such as LLMs, with symbolic solvers enhances performance on sequential decision-making tasks.
 To the best of our knowledge, neurosymbolic techniques have not been explored in the context of graph coloring problems.
-The SOFAI Architecture
-The SOFAI architecture, also referred to as SOFAI-v1 here (Booch et al.
-2021), leverages a rule-based metacognitive control mechanism that dynamically chooses between S1 and S2 solvers based on a confidence threshold associated with S1’s outputs.
+To our knowledge, prior neurosymbolic studies have not targeted the graph coloring decision problem, nor integrated meta-level feedback with a decision-oriented S2; SOFAI-v2 explicitly addresses this gap.
+Conceptually, SOFAI-v2 instantiates a neurosymbolic cognitive architecture in which S1 (neural) handles routine proposals while MC+S2 provide supervisory/controlled processing aligned with Norman–Shallice-style accounts (Norman & Shallice, 1986; Cooper & Shallice, 2000).
+4 The SOFAI Architecture
+The SOFAI architecture, also referred to as SOFAI-v1 here (Booch et al., 2021), leverages a rule-based metacognitive control mechanism that dynamically chooses between S1 and S2 solvers based on a confidence threshold associated with S1’s outputs.
 S1 solvers are experience-based (and usually data-driven) solvers, while S2 solvers are deliberative (and usually symbolic and rule-based) solvers.
-SOFAI-v1 has been instantiated to planning (Fabiano et al.
-2023) and constrained-grid navigation (Ganapini et al.
-2022), showing improvement on the performance compared to either symbolic or data-driven solvers for the same class of problems.
+SOFAI-v1 has been instantiated to planning (Fabiano et al., 2023) and constrained-grid navigation (Ganapini et al., 2022), showing improvement on the performance compared to either symbolic or data-driven solvers for the same class of problems.
 In SOFAI-v1, the metacognitive component only chooses between S1 and S2 solvers, but does not exploit the possible collaboration between these two kinds of solvers, nor does it provide feedback to the solvers if they cannot solve the given problem instance.
-In this paper we consider a generalized version of this architecture, called SOFAI-v2, shown in Figure 2, where the metacognitive component (called MC) can provide feedback and examples to a failing S1 solver (that for this paper is an LLM), calling the S1 solver more than once, until either a correct solution is returned or a maximum number of iterations is reached, which triggers the activation of an S2 solver.
+In this paper we consider a generalized version of this architecture, called SOFAI-v2, shown in Figure 1, where the metacognitive governance (called MC) can provide feedback and examples to a failing S1 solver (that for this paper is an LLM), calling the S1 solver more than once, until either a correct solution is returned or a maximum number of iterations is reached, which triggers the activation of an S2 solver.
+This implements a neurosymbolic control loop: S1 produces routine proposals (automatic), MC performs verification and trend monitoring (supervisory attention), and S2 executes controlled search when conflicts persist, echoing contention scheduling with a supervisory system (Norman & Shallice, 1986; Cooper & Shallice, 2000).
 Formally, let $f_{\text{S1}}(x,M)$ denote the solution produced by S1 for a problem instance $x$ with access to an episodic memory $M$.
 Here, $M$ consists of previously encountered problem-solution pairs generated by SOFAI-v2, i.e., pairs $(x_{j},f_{\text{SOFAI-v2}}(x_{j}))$, where $f_{\text{SOFAI-v2}}(x_{j})$ represents the final solution that could have been generated by either S1 or S2.
+In the context of this paper, $x$ corresponds to a graph coloring decision problem instance defined by a graph $G$ and a color bound $k$.
 When a new problem instance $x$ is presented to S1, a similarity function $\sigma(x,M)$ is used to retrieve a subset $M_{x}\subset M$ of similar past instances, defined as:
@@
 where $x_{j}\in M$ and $\alpha$ is a predefined similarity threshold.
-This subset $M_{x}$ provides S1 with additional contextual information from past solutions, enabling it to leverage the SOFAI-v2’s previous outputs for experience-based adaptation.
-Let $C(f_{\text{S1}}(x,M_{x}))$ denote the confidence score assigned by MC to the solution $f_{\text{S1}}(x,M_{x})$, produced by S1 using both the current problem instance $x$ and the retrieved memory subset $M_{x}$.
-If $C(f_{\text{S1}}(x,M_{x}))\geq\theta_{\text{S1}}$, where $\theta_{\text{S1}}$ is a predefined confidence threshold, the solution is accepted as final.
-Otherwise, MC provides feedback $\mathcal{F}(f_{\text{S1}}(x,M_{x}))$ or generates a simplified example $\mathcal{E}(x)$, where $\mathcal{E}(x)$ is a reduced version of $x$, and calls the S1 solver again with the additional information included in the feedback and the examples.
-MC also monitors the trend in S1’s confidence scores over $k$ iterations to evaluate the improvement.
-Let $C_{i}=C(f_{\text{S1}}(x_{i},M_{x_{i}}))$ represent the confidence score after the $i$-th iteration on similar or modified instances.
-An improvement trend is identified if:
-$\forall i\in\{1,\dots,k-1\},\quad C_{i+1}>C_{i}.$
-If this trend condition is not satisfied after $k$ iterations, MC gives the problem instance to S2, invoking $f_{\text{S2}}(x)$ as a fallback to solve it.
-This ensures robustness, particularly in cases where S1’s adaptation is insufficient to meet confidence requirements, even with feedback, examples, and episodic memory support.
-Experimental Setup and Results
-This section details the methodology and results of experiments designed to evaluate the performance of various solver configurations on graph coloring problems.
-The experiments focus on efficiency and success rate, incorporating a systematic approach to data generation, problem classification, and solver configuration.
-Data Generation and Problem Classification
-Graph coloring problems were generated using the Erdős–Rényi model, allowing for controlled variations in graph complexity through two main parameters:
-Graph Size ($n$): Graphs varied $n\in[5,50]$ vertices, addressing a broad range of problem sizes.
-Edge Formation Probability ($p$): Set with $p\in[0.1,0.9]$, this parameter adjusted the density of edges, thereby influencing the difficulty of the coloring problem.
-Solvability Mix ($m$): The chromatic number for each graph, crucial for defining solvability, was calculated using the DSATUR algorithm.
-Based on this, problems were categorized into different solvability mixes, represented by the tuple $m=(a,b)$, where $a$ is the percentage of solvable problems and $b$ is the percentage of unsolvable problems:
-  • $m=(100,0)$: All problems are solvable within the given constraints.
-  • $m=(50,50)$: A balanced mix with equal solvable and unsolvable scenarios.
-  • $m=(0,100)$: All problems are unsolvable, with the allowed number of colors less than the chromatic number.
-Graphs were saved in DIMACS format to ensure standardization and enhance reproducibility.
-SOFAI-v2 Implementation for Graph Coloring
-The SOFAI-v2 architecture integrates advanced components designed to solve graph coloring problems dynamically.
-Figures 8 and 4 show example prompts provided to LLM (S1) for solving graph coloring.
-Below, we detail the implementation of these components using formal definitions and the functions they perform within the architecture:
-  • Episodic Memory Implementation: The episodic memory, denoted as $M$, stores historical instances of solved graph coloring problems and their solutions.
-This aids the S1 solver by providing access to a set of similar past problems, enhancing the adaptability and effectiveness of the solver:
-    • Memory Structure: Each entry in $M$ consists of graph attributes such as $nodes$, $edges$, $edge\_density$, $chromatic\_number$, and the corresponding $solution$.
-    • Similarity Assessment: A similarity function $\sigma(x,M)$ evaluates the resemblance of the current graph attributes $x$ to those stored in $M$, prioritizing instances where attribute deviations are minimal.
-The similarity function is defined as:
-      $\sigma(x,M)=\{(x_{j},f_{\text{v2}}(x_{j})):x_{j}\in M,\,d(x,x_{j})<\alpha\}$
-      where, $d(x,x\_j)=|attributes(x)-attributes(x\_j)|$,   $\alpha$ threshold.
-  • Improvement Trend Evaluation: If consecutive iterations indicate no improvement in solving capabilities, or if S1 fails to solve a problem within a set limit of five iterations, the S2 solver is invoked to ensure problem resolution.
-  • Example Generation: The Graph Coloring Example Generator, $\mathcal{E}(x)$, simplifies the graph problem space by creating subgraphs.
-This module uses a greedy algorithm to propose potential solutions for these subgraphs, aiding the broader problem-solving strategy by focusing on simpler problem segments.
-  • Feedback Generation and Validator: The Graph Coloring Validator assesses the validity of the proposed solutions and identifies errors such as coloring conflicts or uncolored vertices.
-Feedback denoted as $\mathcal{F}(f_{\text{S1}}(x,M_{x}))$ is provided based on these assessments, critical for the iterative adjustment of the solver’s strategy.
-The feedback process is formally incorporated as part of the metacognitive governance’s evaluation of $C(f_{\text{S1}}(x,M_{x}))$ against a confidence threshold $\theta_{\text{S1}}$.
-This configuration ensures that SOFAI-v2 operates under a robust framework of real-time metacognitive governance, dynamically adapting to the complexity of graph coloring problems through experience-based learning and systematic feedback mechanisms.
-Examples of prompts, feedback types, and generated examples are in Supplementary Material.
+This subset, $M_{x}$, provides S1 with additional contextual information from past solutions, enabling it to leverage SOFAI-v2’s previous outputs for experience-guided context.
+The correctness of solution for the graph coloring decision problem can be checked and evaluated in polynomial time by MC. The correctness score used by MC for graph coloring solutions is:
+$C(f_{S1}(x,M_{x}))=\frac{\sum_{(v,u)\in E}\mathbf{1}[f(v)\neq f(u)]}{|E|}$
+where:
+  • $f(v)$ denotes the assigned color of vertex $v$,
+  • $E$ denotes the set of edges in the graph.
+If this correctness meets or exceeds a threshold $\theta_{\text{S1}}$, solution is accepted.
+Otherwise, MC iteratively provides structured feedback $\mathcal{F}(f_{S1}(x,M_{x}))$ or simplified examples $\mathcal{E}(x)$.
+We set the correctness threshold to 1, since 1 indicates full constraint adherence for the graph-coloring decision task; this ensures S1 outputs are only accepted when a complete, validator-checked coloring is produced.
+MC also monitors S1’s solution correctness across iterative feedbacks, identifying improvement trends defined by:
+$\forall i\in\{1,\dots,m-1\},\quad C_{i+1}>C_{i}$
+Failure to show improvement within $m$ iterations results in MC invoking the fallback S2 solver, ensuring robustness.
+Thus, the iterative feedback mechanism and episodic memory advances SOFAI-v2 over SOFAI-v1, enhancing solver effectiveness via structured feedback and experience-based learning.
+Our theoretical commitments are: (i) dual-solver with metacognitive governance, (ii) episodic memory for similarity-based retrieval, and (iii) verify-and-feedback loops grounded in polynomial-time checks.
+The concrete instantiation (Mistral-7B as S1, DSATUR with backtracking as S2) is an implementation detail; same principles transfer to other domains.
+Detailed algorithm for the architecture is in Supplementary Material Section 8.
+5 Experimental Setup and Results
+This section details methodology and experimental results evaluating various solver configurations on graph coloring problems, emphasizing efficiency and success rate under systematic constraints.
+We also report how iterative feedback and selective escalation can be interpreted as supervisory control improving routine policy performance under increasing constraint conflict (Norman & Shallice, 1986).
+5.1 Data Generation and Problem Classification
+Graph coloring instances were generated using the Erdős–Rényi model, controlling problem complexity through three primary parameters:
+Graph Size ($n$): Varies between $n\in[5,50]$ vertices, encompassing a diverse spectrum of problem sizes.
+Edge Probability ($p$): Adjusted within $p\in[0.1,0.9]$, influencing edge density and problem difficulty.
+Solvability Mix ($m$): The chromatic number ($\chi(G)$) of each generated graph was computed using the DSATUR algorithm.
+Problems were then categorized based on solvability, represented as $m=(a,b)$, where $a$ and $b$ denote percentages of solvable and unsolvable instances respectively:
+  • $m=(100,0)$: All problems solvable within constraints.
+  • $m=(50,50)$: Equal distribution of solvable and unsolvable problems.
+  • $m=(0,100)$: All problems unsolvable; set by choosing the number of available colors $k<\chi(G)$, ensuring infeasibility.
+For each graph size and edge probability pair 100 graph problems were generated and stored in standardized DIMACS format for reproducibility.
+Github Repository: https://github.com/khvedant02/CSP-SOFAI_v2 .
+5.2 SOFAI-v2 Implementation for Graph Coloring
+SOFAI-v2 integrates iterative feedback mechanisms and episodic memory to dynamically resolve graph coloring problems. Figures 7 and 3 illustrate example prompts.
+  • Episodic Memory ($M$): Stores historical graph instances with attributes—vertices, edges, edge density, chromatic number—and solutions.
+A similarity function $\sigma(x,M)$ retrieves relevant past solutions:
+    $M_{x}=\{(x_{j},f_{\text{v2}}(x_{j})):x_{j}\in M,\,d(x,x_{j})<\alpha\}$
+    where $d(x,x_{j})=|attributes(x)-attributes(x_{j})|$, and $\alpha$ is a similarity threshold.
+  • Iterative Feedback Mechanism: If the solver’s correctness score (defined in Section 4) is insufficient, a two-step feedback process occurs:
+    1. 1.
+      Detect Conflict Pairs: Identify adjacent vertices assigned identical colors.
+    2. 2.
+      Template-Based Feedback: Algorithmically generate structured corrections (e.g., “Error: Vertices A and B are adjacent but have the same color”).
+Example of feedbacks is given in Supplementary Material Section 9.
+  • Improvement Evaluation and S2 Invocation: If no solution improvement occurs over five iterations or if correctness remains below threshold, S2 is invoked as a fallback.
+If either solver identifies an instance as unsolvable, the instance is labeled as UNSAT.
+Exceeding the strict 200-second per-instance time limit results in failure, regardless of actual solvability.
+  • Example Generation ($\mathcal{E}(x)$): A greedy-based algorithm generates simpler subgraph instances to guide S1. Examples are detailed further in Supplementary Material Section 9.
+This comprehensive integration ensures SOFAI-v2 leverages both real-time metacognitive governance and experience-guided context for enhanced problem-solving effectiveness.
+Operationally, this mirrors a supervisory system that monitors conflict and recruits controlled processing when routine proposals fail (Norman & Shallice, 1986; Cooper & Shallice, 2000).
 Graph Coloring Problem Prompt New Problem to Solve:
-You are given an undirected graph with 3 colors available. Your task is to assign a color to each vertex such that no two adjacent vertices share the same color.
+You are given an undirected graph with 2 colors available. Your task is to assign a color to each vertex such that no two adjacent vertices share the same color.
 Graph Representation:
@@
 Assign a unique color to each vertex, ensuring that no two vertices connected by an edge have the same color.
-Use no more than 3 distinct colors.
+Use no more than 2 distinct colors.
 Provide the color assignments for each vertex in the format:
@@
 (C 1)
-Please provide the color assignment for the new problem to solve directly below:
-Figure 3: LLM prompt template for $m(100,0)$ for the graph coloring problem without using episodic memory.
-SOFAI-v1 vs SOFAI-v2: SOFAI-v2 introduces significant advancements over its predecessor, SOFAI-v1, by incorporating iterative feedback mechanisms and episodic memory, which enhance the problem-solving capabilities of the S1 solver.
-Unlike SOFAI-v1, where problems not solved by S1 are directly escalated to S2 without iterative refinement, SOFAI-v2 systematically leverages historical data and adaptive feedback to optimize solutions before resorting to S2, thereby improving efficiency and efficacy.
+Please provide the color assignment for the new problem to solve, or respond with "NOT SOLVABLE" if it cannot be solved.
+Figure 2: LLM prompt template for the graph coloring decision problem without using episodic memory.
 Episodic Memory Prompt Problem:
@@
 End of Example
-Figure 4: Example for episodic memory, which is added at the end of LLM prompt template in Fig 8.
-Performance Metrics
-All solver configurations, as detailed in Table 1, were evaluated using two primary metrics: success rate and average time taken.
-The success rate shown in Equation 1 measures the proportion of correct solutions out of the total number of problems, while the average time shown in Equation 2 taken calculates the mean time per problem solved within the time cap.
-The metrics are defined as follows:
+Figure 3: Example for episodic memory, which is added at the end of LLM prompt template in Supplementary Material Section 9.
+SOFAI-v1 vs SOFAI-v2: SOFAI-v2 introduces substantial advancements over previous works of SOFAI (SOFAI-v1), by incorporating iterative feedback mechanisms and episodic memory, which enhance the problem-solving capabilities of the S1 solver.
+Unlike SOFAI-v1, where problems not solved by S1 are directly escalated to S2 without iterative refinement, SOFAI-v2 systematically leverages historical data and adaptive feedback to optimize solutions before resorting to S2, thereby improving efficiency and efficacy.
+Supplementary Material includes exact prompts, feedback templates, and our DSATUR+BT pseudocode.
+5.3 Performance Metrics and Evaluation
+All solver configurations (Table 1) were evaluated under a strict 200-second time limit per instance. Solver performance was measured using:
 $\text{Success Rate}(\%)=\left(\frac{\text{Number of Correct Solutions}}{\text{Total Number of Problems}}\right)\times 100$ (1)
-Average Time $\text{Average Time Taken}=\frac{\sum_{i=1}^{n}t_{i}}{n}$ (2)
-where in Equation 2, $t_{i}$ represents the time taken to solve the $i$-th problem, and $n$ is the total number of problems solved within the time cap.
-Results were aggregated over three randomized trials to ensure reliability and consistency in the findings.
+$\text{Average Time Taken}=\frac{\sum_{i=1}^{n}t_{i}}{n}$ (2)
+where $t_{i}$ is the solving time for the $i$-th successfully solved instance and $n$ the total instances solved within the time limit.
+Instances exceeding this limit are marked as failures, explaining occasional 0% success rates even when solvable theoretically.
+Results reported are averaged over three randomized trials for reliability.
+To better understand solver efficiency beyond strict time constraints, we conducted supplementary experiments without time limits, solely for comparative runtime analysis.
+From a cognitive perspective, increased success with iterative MC reflects benefits of supervisory control over routine responding under higher constraint densities (Norman & Shallice, 1986; Cooper & Shallice, 2000).
+Table 1: Descriptions of solver configurations employed in experiments.
 ID Description
@@
 S2 DSATUR with Backtracking
-SOFAI-v1 Combines S1 and S2 with rule-based metacognition
-SOFAI-v2 Combines S1 and S2 with real-time metacognition
-MC-S1-I[1-5] Iterative metacognitive feedback ranging from one to five iterations to refine S1 solutions.
-Table 1: Descriptions of solver configurations employed in the experiments.
-Results
+SOFAI-v1 Combines S1 and S2 with rule-based MC
+SOFAI-v2 Combines S1 and S2 with iterative MC
+MC-S1-I[1-5] Metacognitive feedback (1–5 iterations) to refine S1 solutions
+5.4 Results
 This section presents an analysis of solver performance by systematically controlling key parameters.
@@
 Tables 2 and 3 provide the success rate and average time taken for each solver configuration under these controlled conditions.
-The Supplementary Material includes results for additional $p$ values to supplement this focused analysis, providing a comprehensive view of solver robustness across varying edge densities.
+The Supplementary Material 11 & 12 includes results for additional $p$ values to supplement this focused analysis, providing a comprehensive view of solver robustness across varying edge densities.
 From these results, we aim to address four key research questions (RQs) that provide deeper insights into SOFAI-v2 performance and robustness.
-RQ 1 What improvement in success rate does SOFAI-v2 demonstrate compared to other solvers?
-Graph Size m(100,0) (in %) m(0,100) (in %) m(50,50) (in %)
+RQ 1 How does SOFAI-v2 success rate compare to that of S1, S2, SOFAI-v1?
+Table 2: Success rates (%) across solvers and problem configurations for graph sizes from 5 to 50 with edge probability $p=0.5$.
+Configurations $m=(100,0)$, $m=(0,100)$, and $m=(50,50)$ represent solvable, unsolvable, and a balanced mix of solvable and unsolvable instances, respectively.
+Graph Size m(100,0) [%] m(0,100) [%] m(50,50) [%]
 S1 S2 SOFAI_v1 SOFAI-v2 S1 S2 SOFAI_v1 SOFAI-v2 S1 S2 SOFAI_v1 SOFAI-v2
@@
 50 0 0 0 0 3.85 0 3.85 53.85 2.11 0 2.11 27.72
-Table 2: Success rates (%) across solvers and problem configurations for graph sizes from 5 to 50 with edge probability $p=0.5$.
-Configurations $m=(100,0)$, $m=(0,100)$, and $m=(50,50)$ represent all solvable, all unsolvable, and a balanced mix of solvable and unsolvable instances, respectively.
-According to the results shown in Table 2, SOFAI-v2, SOFAI-v1, and S2 exhibit comparable performance in fully solvable problem configuration ($m=(100,0)$).
-In contrast, in the unsolvable scenario ($m=(0,100)$) at a graph size of 50, SOFAI-v2 achieves a success rate of 53.85%, which is significantly higher than the 3.85% achieved by both SOFAI-v1 and S1.
-Similarly, in the mixed scenario ($m=(50,50)$), SOFAI-v2 continues to perform well with a 27.72% success rate, far surpassing SOFAI-v1 and S1, each at 2.11%.
-The higher success rates of SOFAI-v2 in challenging problem configurations likely arise from its metacognitive governance that effectively identifies and resolves constraints, particularly in unsolvable configurations.
-This capability allows SOFAI-v2 to adapt more dynamically to complex problem sets, enhancing its performance in environments with a mix of solvable and unsolvable problems.
-Answer to RQ 1: The SOFAI-v2 model shows enhanced performance in tackling unsolvable and mixed problem configurations, surpassing other solvers as detailed in Table 2.
-For example, within the unsolvable scenarios set (with $m=(0,100)$ at $n=50$), SOFAI-v2 achieves a success rate of 53.85%.
-This is approximately a 1298% improvement over the success rates of SOFAI-v1 and Solver S1, which are around 3.85%.
-The calculation for the improvement is derived as follows:
-$\frac{53.85-3.85}{3.85}\times 100\%\approx 1298\%$
-This significant performance boost is attributed to SOFAI-v2’s metacognitive governance system, which optimizes decision-making processes in complex problem configurations.
-RQ 2 How does SOFAI-v2 compare in terms of average time taken for solving problems relative to other solvers?
-Graph Size m(100,0) (in seconds) m(0,100) (in seconds) m(50,50) (in seconds)
-S1 S2 SOFAI_v1 SOFAI-v2 S1 S2 SOFAI_v1 SOFAI-v2 S1 S2 SOFAI_v1 SOFAI-v2
-5 3.23 0.001 3.23 0.92 3.13 0.001 3.13 0.77 3.23 0.001 3.23 0.92
-10 4.56 0.19 4.69 2.21 4.08 0.17 4.14 1.07 5.28 0.22 5.48 3.74
-15 5.61 70.85 55.46 53.23 6.20 72.02 39.89 22.33 5.12 67.08 70.01 85.37
-20 6.79 199.27 154.11 120.86 7.52 198.70 111.00 46.96 6.68 196.22 193.15 195.25
-30 11.53 200.00 137.80 117.10 10.87 200.00 82.30 52.10 9.92 200.00 192.85 193.67
-40 9.69 200.00 153.41 144.54 9.46 200.00 109.71 101.18 10.30 200.00 194.47 193.67
-50 41.98 200.00 200.00 139.17 62.76 200.00 88.93 88.93 21.83 200.00 200.00 194.47
-Table 3: Average time taken (in seconds) for different solvers across graph sizes and problem configurations, highlighting operational efficiency in solving or recognizing unsolvable problems.
-According to the results in Table 3, SOFAI-v2 demonstrates distinct time efficiency advantages across various problem configurations and graph sizes.
-In fully solvable problems ($m=(100,0)$), the average solving time for SOFAI-v2 remains consistent and competitive with other solvers.
-However, in more complex problem configurations, such as all unsolvable problems ($m=(0,100)$) at $n=50$, SOFAI-v2 significantly reduces the average solving time to 88.93 seconds, a 55% decrease compared to S2, requiring up to 200 seconds.
-Similarly, in the mixed scenario ($m=(50,50)$) at the same graph size, SOFAI-v2 further improves efficiency, reducing the solving time to 194.47 seconds, demonstrating a nearly 3% improvement over SOFAI_v1 (200 seconds) and S2 (200 seconds).
-The enhanced time efficiency of SOFAI-v2 in these challenging configurations likely results from its metacognitive governance, which optimizes computational strategies to handle complex and mixed problem configurations more efficiently.
-This capability enables SOFAI-v2 to solve problems faster and manage its computational resources more effectively, particularly in problem configurations where solvability varies widely.
-Answer to RQ 2: In terms of time efficiency, SOFAI-v2 significantly outperforms other solvers, notably in problem configurations involving unsolvable and mixed problems.
-Specifically, at a graph size of 50 within the unsolvable configuration ($m=(0,100)$), the average time taken by SOFAI-v2 to attempt solving these problems is 88.93 seconds, which is approximately 55% less than the 200 seconds typically required by Solver S2.
-The percentage reduction is calculated as follows:
-$\frac{200-88.93}{200}\times 100\%\approx 55\%$
-This enhanced performance is attributed to the metacognitive governance within SOFAI-v2, which accelerates the problem-solving process and increases accuracy in complex problem configurations.
-RQ 3 Does the iterative calling mechanism in SOFAI-v2 enhance the performance of its S1 solver through metacognitive governance?
+Table 3: Average time (seconds) for SOFAI_v1, SOFAI_v2, and S2 across graph sizes and configurations $m(100,0)$, $m(0,100)$, and $m(50,50)$.
+Graph Size $m(100,0)$ (s) $m(0,100)$ (s) $m(50,50)$ (s)
+SOFAI_v1 SOFAI_v2 S2 SOFAI_v1 SOFAI_v2 S2 SOFAI_v1 SOFAI_v2 S2
+5 2 1 0 2 1 0 2 1 0
+10 4 2 0 4 2 0 4 3 0
+15 50 40 51 32 19 59 52 47 54
+20 87 65 82 50 20 97 87 84 107
+30 323 306 320 403 299 400 353 263 350
+40 5043.2 4512.3 4995.8 5006.9 3511.7 4989.4 5012.1 3524.6 4992.7
+50 20087.3 18034.9 19950.5 20044.2 14102.7 19920.6 20051.8 14178.3 19965.0
+Answer to RQ 1: SOFAI-v2 demonstrates substantial improvements in success rates over other solvers, particularly in challenging problem configurations involving unsolvable or mixed solvability instances.
+Table 2 shows that for clearly solvable scenarios ($m=(100,0)$), SOFAI-v2 matches the optimal performance of SOFAI-v1 and S2 across smaller graph sizes.
+However, its key advantage becomes apparent in more complex scenarios:
+  • For unsolvable scenarios ($m=(0,100)$) at graph size 50, SOFAI-v2 achieves a success rate of 53.85%, far surpassing the performance of SOFAI-v1 and S1, which each attain only 3.85%.
+This translates to an improvement of approximately 1298%.
+  • In mixed scenarios ($m=(50,50)$), SOFAI-v2 maintains a success rate of 27.72%, substantially higher compared to just 2.11% by SOFAI-v1 and S1.
+These performance gains arise primarily from SOFAI-v2’s enhanced metacognitive governance, which systematically leverages iterative solver feedback and episodic memory.
+This adaptive capability allows SOFAI-v2 to dynamically resolve constraints and effectively handle complex or initially challenging instances that other solvers cannot reliably solve.
+For the above results we set a fixed time limit of 200 seconds per problem.
+If a solver does not finish within this time, it is counted as a failure.
+The average time taken results for this setting is reported separately in the Supplementary Material Section 10.
+RQ 2 How does SOFAI-v2 compare in terms of average time relative to SOFAI-v1 and S2?
+Answer to RQ 2: SOFAI-v2 demonstrates substantial improvements in time efficiency compared to SOFAI_v1 and S2, especially for complex configurations involving unsolvable or mixed scenarios.
+As detailed in Table 3, in the unsolvable scenario ($m=(0,100)$) at graph size $n=50$, SOFAI-v2 requires approximately 14,000 seconds, whereas both SOFAI_v1 and S2 take around 20,000 seconds.
+This corresponds to a substantial 30% reduction in average solving time.
+Similarly, for mixed solvability scenarios ($m=(50,50)$) at the same graph size ($n=50$), SOFAI-v2 achieves the same notable reduction of 30%, requiring only about 14,000 seconds compared to 20,000 seconds for SOFAI_v1 and S2.
+In fully solvable scenarios ($m=(100,0)$), the improvements remain consistent.
+For instance, at graph size 50, SOFAI-v2 completes solving tasks in roughly 18,000 seconds, a 10% time reduction compared to SOFAI_v1 and S2 (each requiring about 20,000 seconds).
+These time-efficiency gains are attributed to SOFAI-v2’s advanced metacognitive governance, which intelligently manages solver interactions and computational resources based on real-time feedback and memory-driven decision-making.
+This dynamic resource allocation enhances solver efficiency, especially in challenging problem settings.
+To ensure a fair and meaningful comparison of solver speed, we conducted this experiment without imposing any time limit.
+This is appropriate because SOFAI-v1, SOFAI-v2, and S2 are all complete solvers—that is, they can eventually solve or correctly determine the unsolvability of a problem given enough time.
+In contrast, S1 alone does not satisfy this property, so it is excluded from this timing comparison.
+This no-time-limit evaluation allows us to isolate and understand the time-efficiency advantages introduced by metacognitive governance in SOFAI-v2.
+RQ 3 Does increasing the number of iterations in the feedback loop lead to an increased success rate in SOFAI-v2?
 (a) All Solvable ($m=(100,0)$)
@@
 Figure 6: Time efficiency across solvers as a function of edge probability in different problem configurations, showcasing SOFAI-v2’s performance in reducing average time taken.
-The iterative feedback mechanism within SOFAI-v2’s S1 solver shows a clear progression in performance across different graph sizes and problem configurations, as detailed in Figure 5.
-Each iteration, from MC-S1-I1 through MC-S1-I5, demonstrates a quantitative improvement in success rates, indicative of the solver’s enhanced problem-solving capabilities due to metacognitive governance.
-In the fully solvable problems ($m=(100,0)$), the success rates of the initial iteration MC-S1-I1 start at approximately 20% and quickly improve to nearly 80% by MC-S1-I5 for $n=5$.
-This rapid improvement highlights the effectiveness of iterative feedback in optimizing the solver’s approach to solvable problems.
-For the entirely unsolvable problems ($m=(0,100)$), success rates at $n=10$ begin at about 60% during MC-S1-I1 and rise to over 85% by MC-S1-I5.
-This increase of approximately 40% in success rate across iterations underscores the solver’s growing capability to efficiently recognize and handle unsolvable problem configuration without unnecessarily invoking higher-cost computational resources.
-In the balanced mix problem configuration ($m=(50,50)$), iterative feedback leads to substantial improvement across graph sizes.
-Initially, success rates at MC-S1-I1 are around 40% for $n=10$ and increase to about 60% by MC-S1-I5.
-This demonstrates a consistent improvement of around 50% in discerning between solvable and unsolvable problems as iterations progress, illustrating the benefits of metacognitive governance in mixed problem configurations.
-Answer to RQ 3: The iterative calling mechanism of SOFAI-v2 substantially enhances the performance of its S1 solver in various problem configurations.
-For instance, in the configuration where $m=(100,0)$, success rates escalate from 20% in the MC-S1-I1 to nearly 80% by the MC-S1-I5, marking an approximate 300% increase.
-This improvement is calculated as follows:
-$\frac{80-20}{20}\times 100\%=300\%$
-Similarly, in the configuration $m=(0,100)$, the success rates rise by 40%, from 60% in MC-S1-I1 to 85% in MC-S1-I5:
-$\frac{85-60}{60}\times 100\%\approx 40\%$
-In the balanced scenario ($m=(50,50)$), there is a 50% increase in success rates, from 40% in MC-S1-I1 to 60% in MC-S1-I5, demonstrating the solver’s adaptability:
-$\frac{60-40}{40}\times 100\%=50\%$
-The metacognitive governance within SOFAI-v2 facilitates these significant improvements across different configurations, enhancing the solver’s efficiency and adaptability.
-RQ4 How does edge probability ($p$) influence the solvability and time efficiency of graph coloring problems across different solvers?
-The impact of edge probability ($p$) on solvability and time efficiency across solvers was systematically evaluated through success rates and solving times as $p$ ranges from 0.1 to 0.9.
-As $p$ increases, all solvers generally exhibit a decreasing trend in success rates until reaching a plateau, as detailed in Figure 6.
-This trend indicates the increasing complexity of the graph coloring problems with more connections between nodes.
-Regarding time efficiency, there is a noticeable increase in average solving time, peaking before stabilizing, as shown in Figure 7.
-SOFAI-v2 consistently outperforms S1, S2, and SOFAI-v1 in maintaining higher success rates across all edge probabilities, demonstrating enhanced resilience against increasing problem complexity.
-Additionally, SOFAI-v2 consistently requires less time to solve problems across all configurations, highlighting its superior time efficiency.
-In the All Solvable ($m=(100,0)$) configuration, SOFAI-v2 exhibits comparable success rates as the probability $p$ increases, similar to other solvers.
-In the All Unsolvable ($m=(0,100)$) scenario, SOFAI-v2 achieves peak success rates at an edge probability of $p=0.1$, which are 20-40% higher than other solvers (SOFAI-v2 = 94%, SOFAI-v1 = 74%, Solver S1 = 68%, Solver S2 = 51%).
-This advantage is maintained with an approximate 10-20% lead as $p$ increases, as illustrated in Figure 6.
-Additionally, in this configuration, SOFAI-v2 solves problems 60-80% faster than its counterparts at mid-range probabilities ($p=0.1$ to $0.5$).
-For instance, at $p=0.1$, the solving times are as follows: SOFAI-v2 = 20 seconds, SOFAI-v1 = 55 seconds, and Solver S2 = 112 seconds, as depicted in Figure 7.
-In the Balanced Mix ($m=(50,50)$) scenario, the success rate for SOFAI-v2 peaks at approximately 72% at $p=0.1$, which is about 15-25% higher than the rates observed for SOFAI-v1, Solver S1, and Solver S2.
-The time efficiency in this scenario also notably surpasses other solvers, with reductions in solving times of 40-50%.
-Specifically, at $p=0.1$, the times are: SOFAI-v2 = 70 seconds, SOFAI-v1 = 100 seconds, and Solver S2 = 108 seconds.
-Answer to RQ 4: Edge probability ($p$) substantially influences solvability and time efficiency, with SOFAI-v2 showing superior adaptability and operational efficiency.
-In problem configurations with increasing complexity, SOFAI-v2 sustains higher success rates and achieves significant time efficiency compared to other solvers, affirming its effective metacognitive governance.
-Discussion
-This study presented SOFAI-v2, a neurosymbolic architecture that integrates a fast, experience-based S1 powered by an LLM with a slow, deliberative S2 utilizing the DSATUR graph coloring algorithm governed by real-time metacognitive governance.
-Our experiments evaluated the performance of SOFAI-v2 on graph coloring problems with sizes ranging from 5 to 50, generated using the Erdős–Rényi model with an edge probability of $p=0.5$.
-We assessed the solvers across three problem configurations: all solvable ($m=(100,0)$), all unsolvable ($m=(0,100)$), and a balanced mix ($m=(50,50)$).
-The discussion addresses the research questions posed.
-Integrating metacognitive governance and episodic memory in SOFAI-v2 demonstrates a notable enhancement in solving graph coloring problems, significantly improving success rates and reducing computational times across various graph sizes and problem complexities.
-Success rates in complex scenarios like the all-unsolvable configuration reveal SOFAI-v2’s superior adaptability, achieving up to a 53.85% success rate for larger graphs, a substantial improvement over SOFAI-v1 and traditional symbolic solvers.
-Additionally, iterative metacognitive feedback in SOFAI-v2’s System 1 has proven effective in refining problem-solving strategies, increasing efficiency, and enhancing the solver’s ability to differentiate between solvable and unsolvable problems.
-These capabilities underscore the value of combining neural and symbolic approaches to overcome limitations in traditional CSP solvers, demonstrating SOFAI-v2’s potential in dynamic and demanding CSP environments.
-Implications and Comparisons with Prior Work
+Answer to RQ 3: Yes, the iterative calling mechanism substantially enhances the performance of SOFAI-v2’s S1 solver, driven by effective metacognitive governance.
+As illustrated in Figure 4, repeated solver invocations with structured feedback consistently improve success rates across all problem configurations:
+  • In fully solvable problems ($m=(100,0)$), the initial success rate at iteration MC-S1-I1 (around 20% for graph size $n=5$) improves substantially to nearly 80% by iteration MC-S1-I5, a notable increase of 300%.
+  • For completely unsolvable instances ($m=(0,100)$) at $n=10$, success rates begin at approximately 60% (MC-S1-I1) and increase to 85% by MC-S1-I5, an improvement of 40%.
+  • In balanced solvability scenarios ($m=(50,50)$) at graph size $n=10$, success rates also see meaningful improvements, increasing from approximately 40% (MC-S1-I1) to 60% (MC-S1-I5), representing a 50% improvement.
+These consistent improvements underscore how SOFAI-v2’s metacognitive governance improves solver responses through iterative feedback, enhancing overall adaptability and efficiency across diverse problem settings.
+For the above results, we set a fixed time limit of 200 seconds per problem.
+If a solver does not finish within this time, it is counted as a failure.
+RQ4 How does the density of graph influence success rate and time across the considered solvers?
+Answer to RQ 4: Edge probability ($p$) substantially affects both solvability and solving time across graph coloring solvers, with SOFAI-v2 demonstrating superior adaptability and efficiency.
+Figure 5 shows that increasing $p$ generally reduces success rates due to higher complexity from additional edges.
+However, SOFAI-v2 consistently maintains higher success rates compared to other solvers:
+  • In fully solvable problems ($m=(100,0)$), all solvers maintain high success rates as $p$ increases, with SOFAI-v2 performing comparably to S2 and SOFAI-v1.
+  • In completely unsolvable scenarios ($m=(0,100)$), SOFAI-v2 notably outperforms others, particularly at lower edge probabilities.
+For instance, at $p=0.1$, SOFAI-v2 achieves a 94% success rate, considerably surpassing SOFAI-v1 (74%), S1 (68%), and S2 (51%).
+This advantage remains pronounced (10-20%) even as $p$ increases.
+  • In mixed scenarios ($m=(50,50)$), SOFAI-v2 again excels, achieving a peak success rate of about 72% at $p=0.1$, approximately 15-25% higher than other solvers.
+Regarding time efficiency (Figure 6), as $p$ increases, solving times rise initially and then stabilize. However, SOFAI-v2 consistently requires less time across all problem configurations:
+  • In the unsolvable scenario ($m=(0,100)$), SOFAI-v2 solves problems substantially faster at lower and mid-range probabilities.
+For instance, at $p=0.1$, SOFAI-v2 averages 20 seconds per problem, far less than SOFAI-v1 (55 seconds) and S2 (112 seconds).
+  • For mixed solvability scenarios ($m=(50,50)$), SOFAI-v2 again demonstrates superior efficiency, 40-50% faster.
+Specifically, at $p=0.1$, SOFAI-v2 solves problems in about 70 seconds, compared to SOFAI-v1’s 100 seconds and S2’s 108 seconds.
+Overall, SOFAI-v2’s robust metacognitive governance enables it to adapt efficiently and effectively to varying complexities induced by changing edge probabilities, achieving higher success rates and superior time efficiency.
+For the above results we set a fixed time limit of 200 seconds per problem.
+If a solver does not finish within this time, it is counted as a failure.
+5.5 Other Considerations
+We also explicitly address concerns regarding context window limitations inherent in modern LLMs.
+Although iterative prompting in our method increases token consumption, our analyses confirm that token usage remains comfortably within the Mistral-7B model’s 32k-token context limit, even under extreme conditions (graph size 100, edge probability 0.9, and five iterations):
+  • Graph Representation (DIMACS format): approximately 1500 tokens
+  • Initial Prompt & Instructions: approximately 500 tokens
+  • Metacognitive Feedback per iteration (×5): approximately 2500 tokens (500 tokens each)
+  • Episodic Memory Retrieval per iteration (×5): approximately 5000 tokens (1000 tokens each)
+  • LLM Response per iteration (×5): approximately 5000 tokens (1000 tokens each)
+  • Accumulated Prompt per iteration (including previous responses, ×5): approximately 15,000 tokens (3000 tokens each iteration)
+In total, even the worst-case scenario results in approximately 24,500 tokens, well within Mistral-7B’s 32k-token context window.
+Consequently, context-window overflow is not an issue in our experiments.
+Operationally, the metacognitive governance budgets information (problem, feedback, episodic examples) so that fast/automatic proposals remain verifiable within resource bounds, aligning with supervised control over routine processing (Norman & Shallice, 1986; Cooper & Shallice, 2000).
+Additionally, to account for the inherent variability in LLM output across runs, we repeat all experiments involving SOFAI-v1, SOFAI-v2, and S1 three times using the same set of problem instances.
+For success rate evaluation, a problem is considered a success if the solver succeeds in any of the three runs.
+For average time taken, we report the minimum time observed across the three trials.
+This choice reflects a conservative estimation of each system’s upper potential in ideal conditions and avoids penalizing performance due to stochastic failures.
+Such a best-of-$n$ (here n=3n{=}3) strategy is commonly used in LLM evaluations to mitigate randomness and highlight consistent solver capabilities (Kang et al., 2025; Chow et al., 2024).
+Cognitively, this also controls for noise in routine responding while assessing whether supervisory interventions reliably reduce errors.
+We also note that SOFAI-based solvers (and S1) can declare UNSAT without a constructive certificate, whereas S2’s UNSAT arises from explicit search.
+This asymmetry can bias results on mixes containing unsatisfiable instances.
+However, the systematic variation in success rates across graph sizes and densities indicates the behavior is not arbitrary and reflects nontrivial problem sensitivity.
+A full proof-producing pipeline for UNSAT is orthogonal to our current scope; for automatic and quick evaluation, we accept such declarations, and we flag this as a limitation to be addressed in future work
+5.6 Implications and Comparisons with Prior Work
 This study broadens the scope of neurosymbolic systems by applying them to graph coloring problems.
-It demonstrates that integrating LLMs with traditional symbolic solvers, guided by metacognitive governance, significantly enhances solutions’ accuracy and efficiency.
-Unlike previous applications that primarily leveraged LLMs for sequential decision-making tasks (Valmeekam et al.
-2022; Stechly, Valmeekam, and Kambhampati 2024), SOFAI-v2 directly addresses their constraint adherence limitations through iterative feedback and strategic use of symbolic reasoning.
-This approach aligns with cognitive decision-making theories and shows a better success rate and time efficiency, illustrating the practical benefits of hybrid architectures in complex constraint satisfaction problems.
-Conclusion
+It demonstrates that integrating LLMs with traditional symbolic solvers, guided by metacognitive governance, enhances accuracy and efficiency.
+Unlike previous applications that primarily leveraged LLMs for sequential decision-making tasks (Valmeekam et al., 2022; Stechly et al., 2024), SOFAI-v2 directly addresses constraint-adherence limitations through iterative feedback and strategic use of symbolic reasoning.
+Viewed through cognitive control, S1 provides routine proposals, while MC and S2 supply supervisory/controlled processing when conflicts persist, consistent with contention scheduling and supervisory attention (Norman & Shallice, 1986; Cooper & Shallice, 2000).
+This neurosymbolic framing clarifies why SOFAI-v2 improves success rate and time efficiency on constraint satisfaction problems.
+6 Conclusion and Future Work
 We introduced SOFAI-v2, a neurosymbolic architecture that integrates a fast, episodic memory-based LLM (S1) with a slow, deliberative symbolic solver (S2), enhanced by real-time metacognitive governance.
@@
 SOFAI-v2 employs iterative metacognitive feedback and episodic memory-based iterations to enable S1 to refine its response.
-This effectively overcomes the limitations of LLMs in constraint adherance of CSPs.
-By leveraging episodic memory and adaptive feedback, SOFAI-v2 significantly enhances the accuracy and efficiency of solving CSPs, demonstrating the potential of cognitive-inspired AI architectures in managing complex challenges.
-Future work will scale SOFAI-v2 to tackle larger and more varied problem sets, explore different graph models, and apply the architecture to other types of CSPs and sequential decision-making tasks.
-Enhancing the metacognitive components with advanced learning algorithms and incorporating alternative neural models are anticipated to boost its adaptability and efficiency further, contributing to developing robust, scalable solutions for intricate AI applications.
-Appendix A Prompt, Example, and Feedback Examples
+This effectively overcomes the limitations of LLMs in constraint adherence for CSPs.
+By leveraging episodic memory and adaptive feedback, SOFAI-v2 improves the accuracy and efficiency of solving CSPs, indicating the value of cognitively informed, neurosymbolic architectures.
+The observed improvements are consistent with dual-process theories: S1 generates a fast, potential solution; the metacognitive controller checks it for constraint adherence and uncertainty, requests targeted revisions when needed, and escalates to S2 only when necessary; S2 then performs slow, symbolic search as invoked.
+Episodic retrieval provides additional context from prior instances to support S1’s proposal and the controller’s checks.
+In this arrangement, SOFAI-v2 realizes a control loop that finds a trade-off between speed and reliability in a resource-constrained setting.
+Future work will scale SOFAI-v2 to larger and more varied problem sets, explore different graph models, and apply the architecture to other CSPs and sequential decision-making tasks.
+We will also investigate learning-based MC policies and richer memory retrieval, and examine correspondences to cognitive-control signals (e.g., conflict costs) that could enable tighter links to computational cognitive neuroscience.
+Concretely, we will design human-aligned probes (e.g., speed–accuracy manipulations, conflict-cost perturbations, and memory cueing/interference) to test whether SOFAI-v2’s control dynamics mirror classic findings on supervisory attention and conflict-driven effort allocation.
+Appendix
+7 Modified DSATUR Algorithm with Backtracking
+Given the graph coloring problem $(G,C,k)$, with an undirected graph $G=(V,E)$, a set of available colors $C={c_{1},c_{2},\dots,c_{k}}$, and the maximum allowed number of colors $k$, the modified DSATUR algorithm systematically explores assignments using backtracking.
+Below is a detailed, mathematically rigorous description:
+Definitions:
+  • Saturation degree $d_{sat}(v)$ for a vertex $v\in V$ is the cardinality of the set of distinct colors assigned to its adjacent vertices:
+  • Vertex degree $d(v)$ is the number of edges incident to $v$.
+Algorithm Steps:
+  1. 1.
+    Initialization: Assign an initial coloring $f(v)=\emptyset$ for all $v\in V$ (all vertices uncolored).
+  2. 2.
+    Compute Saturation Degree: For each uncolored vertex $v$, compute its saturation degree $d_{sat}(v)$.
+  3. 3.
+    Select Vertex: Identify vertex $v^{*}$ that maximizes $d_{sat}(v)$. If multiple vertices have equal highest saturation degree, select the one with the highest vertex degree $d(v)$.
+  4. 4.
+    Color Assignment: Assign to $v^{*}$ the smallest color $c\in C$ that satisfies the condition:
+    $f(v^{*})\neq f(u),\quad\forall u\text{ adjacent to }v^{*}$
+  5. 5.
+    Backtracking: If no valid color can be assigned, recursively backtrack:
+    • Undo the previous assignment and select the next smallest available color.
+    • Continue this recursive search until a valid color assignment is found or all possibilities are exhausted.
+  6. 6.
+    Termination: Repeat steps 2-5 until all vertices are successfully colored, or conclusively determine no valid coloring exists under the provided constraint of $k$ colors.
+This rigorous exploration via backtracking guarantees the determination of a feasible coloring if one exists within the constraint of using no more than $k$ colors.
+8 Algorithm SOFAI-v2
+Algorithm Description.
+Given a graph $G=(V,E)$ and a color bound $k$, SOFAI-v2 runs an S1–MC loop for at most $T$ iterations.
+In each iteration, S1 is prompted with the graph, $k$, and the accumulated history.
+S1 proposes a coloring assignment, which is checked for feasibility.
+If valid, the algorithm returns the assignment.
+If invalid, MC extracts the conflicts, generates minimal checkable feedback, and appends this to the history and trend signals.
+If repeated patterns or the iteration limit is reached, MC escalates to S2.
+S2 runs a DSATUR backtracking solver for the fixed-$k$ decision problem.
+If a valid coloring is found, it is returned and stored in memory; otherwise, the algorithm reports “NOT SOLVABLE”.
+Algorithm 1 SOFAI-v2 for Graph Coloring (Decision, fixed $k$)
+1:Undirected graph $G=(V,E)$, color limit $k$, max S1 iterations $T$, memory $\mathcal{M}$
+2:Coloring assignment or “NOT SOLVABLE”
+[…]
+[…]
+[…]
+6:for $t=1$ to $T$ do
+7:  $prompt\leftarrow\text{GenerateGCPrompt}(G,k,seed,history)$
+8:  $resp\leftarrow\text{CallS1}(prompt)$
+9:  $assign\leftarrow\text{ParseColoring}(resp)$
+10:  $(ok,viol)\leftarrow\text{CheckFeasibility}(G,k,assign)$
+11:  if $ok=\text{true}$ then
+12:   return assignassign
+13:  end if
+14:  $C\leftarrow\text{ExtractConflicts}(viol)$
+15:  $H\leftarrow\text{InducedSubgraph}(G,C)$
+16:  $f\leftarrow\text{FormatFeedback}(H,C,assign)$
+17:  $history\leftarrow history\cup\{(assign,f)\}$
+18:  $trends\leftarrow trends\cup\{\text{FeedbackSignature}(f)\}$
+19:  if $\text{EscalationRule}(trends,t)$ then
+20:   break $\triangleright$ escalate to S2
+21:  end if
+22:  $seed\leftarrow\text{UpdateSeed}(assign,f)$
+23:end for
+24:
+25:// S2 fallback (decision, fixed $k$)
+[…]
+27:if $status=\text{SAT}$ then
+28:  $\text{UpdateMemory}(\mathcal{M},G,k,assign)$
+29:  return assignassign
+30:else
+31:  return “NOT SOLVABLE”
+32:end if
+9 Prompt, Example, and Feedback Examples
 This section discusses prompt type, feedback types, and generated examples.
@@
 (C 1)
-Please provide the color assignment for the new problem to solve, or respond with ”NOT SOLVABLE” if it cannot be solved.
-Figure 8: LLM prompt template for $m(0,100)$ and $m(50,50)$ for the graph coloring problem without using episodic memory.
+Please provide the color assignment for the new problem to solve, or respond with "NOT SOLVABLE" if it cannot be solved.
+Figure 7: LLM prompt template for the graph coloring decision problem without using episodic memory.
 Graph Coloring Problem Prompt with Episodic Memory Example New Problem to Solve:
@@
 This example is provided to demonstrate a previously successful coloring strategy for a similar subgraph, which may be helpful in solving the current problem.
-Please provide the color assignment for the new problem to solve directly below, or respond with ”NOT SOLVABLE” if it cannot be solved.
+Please provide the color assignment for the new problem to solve directly below, or respond with "NOT SOLVABLE" if it cannot be solved.
 Figure 8: LLM prompt template for $m(0,100)$ and $m(50,50)$ for the graph coloring problem, incorporating an episodic memory example to aid problem-solving.
@@
 Figure 11: Example of a correctly colored subgraph is generated as part of the feedback to illustrate the correct coloring assignment for less number of nodes.
-Appendix B Results - Success Rate
+10 Results - Average time Solver Comparison
+Table 4 reports the average time taken by different solver configurations (S1, S2, SOFAI_v1, and SOFAI-v2) under a fixed time limit of 200 seconds.
+These results are obtained from the same experiment used to evaluate success rates in Section 5.4.
+Instances not solved within the allotted 200 seconds are excluded from the average time calculation.
+Cells with a dash (--) indicate that the solver failed to solve any instance in that configuration within the time limit, resulting in no valid average time.
+These values help assess not only whether solvers succeed but also how quickly they do so under realistic constraints.
+Table 4: Average time taken (in seconds) by solvers under a 200-second time limit.
+Results are reported for each graph size and solvability configuration.
+Missing values (dashes) indicate that no problem was solved within the time limit for that solver and configuration.
+These results correspond to the same experiments used in the success rate analysis in RQ1-Section 5.4.
+Graph Size $m(100,0)$ (s) $m(0,100)$ (s) $m(50,50)$ (s)
+S1 S2 SOFAI_v1 SOFAI-v2 S1 S2 SOFAI_v1 SOFAI-v2 S1 S2 SOFAI_v1 SOFAI-v2
+5 3.23 0.001 3.23 3.92 3.13 0.001 3.13 3.77 3.25 0.001 3.23 3.92
+10 3.60 0.19 3.75 3.21 3.80 0.17 3.90 3.07 3.54 0.22 3.85 3.74
+15 3.80 54.25 32.90 31.21 3.85 60.36 34.95 28.21 3.21 56.32 53.92 33.92
+20 3.85 83.22 54.95 48.72 3.90 99.06 55.28 42.31 3.87 108.9 71.98 55.63
+30 - - - - 3.92 - 3.94 19.92 3.92 - 3.97 24.23
+40 - - - - 3.95 - 3.96 22.94 3.95 - 3.99 23.25
+50 - - - - 3.96 - 3.97 21.76 3.97 - 3.97 24.45
+11 Results - Success Rate
 This section explores the success rates of different solvers in solving graph coloring problems across different problem configurations ($m$).
@@
 These graphs show the robustness of SOFAI-v2 in terms of success rate compared to other solvers.
+For the above results we set a fixed time limit of 200 seconds per problem. If a solver does not finish within this time, it is counted as a failure.
 Figure 12: Success Rate (SR) of different solvers vs Graph size ($n$) across edge probabilities ($p$) for problem configuration ($m=(50,50)$)
@@
 Figure 14: Success Rate (SR) of different solvers vs Graph size ($n$) across edge probabilities ($p$) for problem configuration ($m=(100,0)$)
-Appendix C Results - Average Time Take
-This section explores the time taken by different solvers to solve graph coloring problems across different problem configurations ($m$).
+12 Results - Average Time Taken
+This section explores the time different solvers take to solve graph coloring problems across different problem configurations ($m$).
 Figure 15, Figure 16, and Figure 17 demonstrate how each solver—System 1, System 2, SOFAI-v1, and SOFAI-v2—performs across varying edge probabilities ($p$).
-These graphs show the time efficiency of SOFAI-v2 compared to System 2 and SOFAI-v1.
+These graphs show the superiority in time efficiency of SOFAI-v2 compared to System 2 and SOFAI-v1.
+For the above results we set a fixed time limit of 200 seconds per problem. If a solver does not finish within this time, it is counted as a failure.
 Figure 15: Average time of different solvers vs Graph size ($n$) across edge probabilities ($p$) for problem configuration ($m=(50,50)$)